### v1

# 3DGS-LM: Faster Gaussian-Splatting Optimization with Levenberg-Marquardt

Abstract
We present 3DGS-LM, a new method that accelerates the reconstruction of 3D Gaussian Splatting (3DGS) by replacing its ADAM optimizer with a tailored Levenberg-Marquardt (LM). Existing methods reduce the optimization time by decreasing the number of Gaussians or by improving the implementation of the differentiable rasterizer. However, they still rely on the ADAM optimizer to fit Gaussian parameters of a scene in thousands of iterations, which can take up to an hour. To this end, we change the optimizer to LM that runs in conjunction with the 3DGS differentiable rasterizer. For efficient GPU parallization, we propose a caching data structure for intermediate gradients that allows us to efficiently calculate Jacobian-vector products in custom CUDA kernels. In every LM iteration, we calculate update directions from multiple image subsets using these kernels and combine them in a weighted mean. Overall, our method is 30% faster than the original 3DGS while obtaining the same reconstruction quality. Our optimization is also agnostic to other methods that acclerate 3DGS, thus enabling even faster speedups compared to vanilla 3DGS.
Refer to caption Refer to caption
Figure 1: Our method accelerates 3D Gaussian Splatting (3DGS) [20] reconstruction by replacing the ADAM optimizer with a tailored Levenberg-Marquardt. Left: starting from the same initialization, our method converges 30% faster on the Tanks&Temples “Train” scene. Right: after the same amount of time, our method produces higher quality renderings (e.g., better brightness and contrast).
1 Introduction
Novel View Synthesis (NVS) is the task of rendering a scene from new viewpoints, given a set of images as input. NVS can be employed in Virtual Reality applications to achieve photo-realistic immersion and to freely explore captured scenes. To facilitate this, different 3D scene representations have been developed [28, 20, 30, 3, 36, 2]. Among those, 3DGS [20] (3D Gaussian-Splatting) is a point-based representation that parameterizes the scene as a set of 3D Gaussians. It offers real-time rendering and high-quality image synthesis, while being optimized from a set of posed images through a differentiable rasterizer.
3DGS is optimized from a set of posed input images that densely capture the scene. The optimization can take up to an hour to converge on high-resolution real-world scene datasets with a lot of images [42]. It is desirable to reduce the optimization runtime which enables faster usage of the reconstruction for downstream applications. Existing methods reduce this runtime by improving the optimization along different axes. First, 3DGS renders images with a tile-based, differentiable rasterizer that is implemented in CUDA. In every iteration, the Gaussians get updated with gradient descent by backpropagating a rendering loss through the rasterizer. By improving the speed of the forward- and backward-pass, recent methods accelerate the optimization [11, 27, 41, 14]. Second, in 3DGS the number of Gaussians is gradually grown during optimization, which is known as densification. To facilitate this, 3DGS accumulates positional gradients over multiple iterations and uses these statistics to split and prune Gaussians. Recently, GS-MCMC [22], Taming-3DGS [27], Mini-Splatting [13], and Revising-3DGS [5] propose novel densification schemes that reduce the number of required Gaussians to represent the scene. This makes the optimization more stable and also faster, since fewer Gaussians must be optimized.
Despite these improvements, the optimization still takes significant resources, requiring thousands of gradient descent iterations to converge. To this end, we aim to reduce the runtime by improving the underlying optimization during 3DGS reconstruction. More specifically, we propose to replace the widely used ADAM [23] optimizer with a tailored Levenberg-Marquardt (LM) [29]. This allows us to accelerate 3DGS reconstruction (Fig. 1 left) by over 30% on average. Concretely, we propose a highly-efficient GPU parallization scheme for the preconditioned conjugate gradient (PCG) algorithm within the inner LM loop in order to obtain the respective update directions. To this end, we extend the differentiable 3DGS rasterizer with custom CUDA kernels that compute Jacobian-vector products. Our proposed caching data structure for intermediate gradients (Fig. 3) then allows us to perform these calculations fast and efficiently in a data-parallel fashion. In order to scale caching to high-resolution image datasets, we calculate update directions from multiple image subsets and combine them in a weighted mean. Overall, this allows us to improve reconstruction time by 30% compared to state-of-the-art 3DGS baselines while achiving the same reconstruction quality (Fig. 1 right).
To summarize, our contributions are:
  • we propose a tailored 3DGS optimization based on Levenberg-Marquard that improves reconstruction time by 30% and which is agnostic to other 3DGS accleration methods.
  • we propose a highly-efficient GPU parallization scheme for the PCG algorithm for 3DGS in custom CUDA kernels with a caching data structure to facilitate efficient Jacobian-vector products.
2 Related Work
2.1 Novel-View-Synthesis
Novel-View-Synthesis is widely explored in recent years [18, 2, 36, 28, 20, 3, 30]. Neural Radiance Fields (NeRF) [28] are particularly successful, because they achieve highly photo-realistic image synthesis results by optimizing MLP weights through volume rendering. To improve training time from days to minutes, NeRF was combined with explicit representations like voxel grids [15, 35], hash grids [30], or points [40].
3D Gaussian Splatting (3DGS) [20] extends this idea by representing the scene as a set of 3D Gaussians, that are rasterized into 2D splats and then $\alpha$-blended into pixel colors. The approach gained popularity, due to the real-time rendering capabilities of high quality images. Since its inception, 3DGS was improved along several axes (next to improving optimization runtime). Recent methods improve the quality of rendered images [43, 17, 26] and the efficiency during rendering [31, 34]. Others obtain better surface reconstructions [19, 16], reduce the memory requirements of the Gaussians [32], and enable training and rendering of large-scale scenes [21, 46]. We similarly adopt 3DGS as our scene representation, but focus on improving the per-scene optimization time.
2.2 Speed-Up Gaussian Splatting Optimization
Obtaining a 3DGS scene reconstruction can be accelerated in several ways. One line of work reduces the number of Gaussians by changing the densification heuristics [22, 27, 13, 5, 26]. Recent methods focusing on sparse-view reconstruction train a neural network as data-driven prior, that directly outputs Gaussians in a single forward pass [12, 25, 8, 7, 6]. In contrast, we focus on the dense-view and per-scene optimization setting, i.e., we do not require a data prior. Other works improve the optimization runtime by improving the implementation of the underlying differentiable rasterizer [41, 11, 27, 14]. We demonstrate that our method is compatible with these approaches, i.e., our optimizer can be plugged into these methods to even further accelerate the optimization.
2.3 Gauss-Newton Optimization For 3D Reconstruction Tasks
NeRF and 3DGS are typically optimized with stochastic gradient descent (SGD) optimizers like ADAM [23] for thousands of iterations. In contrast, many works in RGB-D fusion employ the Gauss-Newton (or Levenberg-Marquardt) algorithms to optimize objectives for 3D reconstruction tasks [47, 48, 38, 37, 9, 10]. By doing so, these methods can quickly converge in an order of magnitude fewer iterations than SGD. Motivated by this, we aim to accelerate 3DGS optimization by adopting the Levenberg-Marquardt algorithm as our optimizer. Rasmuson et al. [33] implemented the Gauss-Newton algorithm for reconstructing a NeRF based on voxel grids. Their technical approach is related to ours, but we implement it for the 3DGS representation in a different way. Concretely, we subsample images in every iteration and introduce a caching data structure. This allows us to achieve state-of-the-art rendering quality, while significantly accelerating the optimization in comparison to existing methods.
3 Method
Figure 2: Method Overview. We accelerate 3DGS optimization by framing it in two stages. First, we use the original ADAM optimizer and densification scheme to arrive at an initialization for all Gaussians. Second, we employ the Levenberg-Marquardt algorithm to finish optimization.
Our pipeline is visualized in  Fig. 2. First, we obtain an initialization of the Gaussians from a set of posed images and their SfM point cloud as input by running the standard 3DGS optimization (Sec. 3.1). In this stage the Gaussians are densified, but remain unconverged. Afterwards, we finish the optimization with our novel optimizer. Concretely, we optimize the sum of squares objective with the Levenberg-Marquardt (LM) [29] algorithm (Sec. 3.2), which we implement in efficient CUDA kernels (Sec. 3.3). This two-stage approach accelerates the optimization compared to only using first-order optimizers.
3.1 Review of Gaussian-Splatting
3D Gaussian Splatting (3DGS)  [20] models a scene as a set of 3D Gaussians, each of which is parameterized by a position, rotation, scaling, and opacity. The view-dependent color is modeled by Spherical Harmonics coefficients of order 3. To render an image of the scene from a specific viewpoint, all visible Gaussians are first projected into 2D splats with a tile-based differentiable rasterizer. Afterwards, the splats are $\alpha$-blended per-pixel to obtain a final pixel color. To fit the Gaussians to image observations, a rendering loss $\mathcal{L}$ is optimized with the ADAM [23] optimizer w.r.t. all Gaussian parameters $\mathbf{x}$:
$\mathcal{L}(\mathbf{x}){=}\frac{1}{N}\sum_{i=1}^{N}(\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i}){+}\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i})))$ (1)
where $\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.8$, $\hat{c}_{i}$ is the rendered color, and $c_{i}$ the ground-truth color for one color channel of one pixel. Typically, ADAM uses a batch size of 1, meaning a random image per iteration is sampled to perform an update step. The number of Gaussians are initialized from the SfM points and is gradually grown during the first half of the optimization, which is known as densification [20].
3.2 Levenberg-Marquardt Optimization for 3DGS
We employ the LM algorithm for optimization of the Gaussians by reformulating the rendering loss as a sum of squares energy function:
$E(\mathbf{x}){=}\sum_{i=1}^{N}(\sqrt{\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i})}^{2}{+}\sqrt{\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i}))}^{2})$ (2)
where we have in total $N{=}6HWP$ residuals for $P$ images with $H{x}W$ pixels each and 3 color channels. We take the square root of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses, to convert Eq. 1 into the required form for the LM algorithm. In other words, we use the identical objective, but a different optimizer. In contrast to ADAM, the LM algorithm requires a large batch size (ideally all images) for every update step to achieve stable convergence [29]. In practice, we select large enough subsets of all images to ensure reliable update steps (see Sec. 3.3.2 for more details).
3DGS uses the structural similarity index measure (SSIM) as loss function during optimization. In SSIM, the local neighborhood of every pixel gets convolved with Gaussian kernels to obtain the final per-pixel score [39]. As a consequence, gradients from the loss flow from every pixel to all Gaussians in the local neighborhood of that pixel. In contrast, the $\mathcal{L}_{1}$ loss only provides gradients from every pixel to the Gaussians along the corresponding ray. For simplicity, we approximate the gradient flow in $\mathcal{L}_{SSIM}$ by backpropagating the per-pixel scores only to the center pixels (ignoring the contribution to other pixels in the local neighborhood). We implement it following the derivation of Zhao et al. [45]. This approximation allows us to keep rays independent from each other when calculating the Jacobian, which is a desirable property for our CUDA kernel implementation (see Sec. 3.3 for more details).
Obtaining Update Directions
In every iteration of our optimization we obtain the update direction $\Delta\in\mathbb{R}^{M}$ for all $M$ Gaussian parameters by solving the normal equations:
$(\mathbf{J}^{T}\mathbf{J}+\lambda_{reg}\textit{diag}(\mathbf{J}^{T}\mathbf{J}))\Delta=\mathbf{J}^{T}\mathbf{F}(\mathbf{x})$ (3)
where $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{N}$ is the residual vector corresponding to Eq. 2 and $\mathbf{J}\in\mathbb{R}^{NxM}$ the corresponding Jacobian matrix.
In a typical dense capture setup, we optimize over millions of Gaussians and have hundreds of high-resolution images. Even though $\mathbf{J}$ is a sparse matrix (each row only contains non-zero values for the Gaussians that contribute to the color of that pixel), it is therefore not possible to materialize $\mathbf{J}$ in memory. Instead, we employ the preconditioned conjugate gradient (PCG) algorithm, to solve Eq. 3 in a matrix-free fashion. We implement PCG in custom CUDA kernels, see Sec. 3.3 for more details.
Apply Parameter Update
After we obtained the solution $\Delta$, we run a line search to find the best scaling factor $\gamma$ for updating the Gaussian parameters:
$\min_{\gamma}E(\mathbf{x}_{k}-\gamma\Delta)$ (4)
In practice, we run the line search on a 30% subset of all images, which is enough to get a reasonable estimate for $\gamma$, but requires fewer rendering passes. Afterwards, we update the Gaussian parameters as: $\mathbf{x}_{k+1}{=}\mathbf{x}_{k}-\gamma\Delta$. Similar to the implementation of LM in CERES [1], we adjust the regularization strength $\lambda_{reg}$ after every iteration based on the quality of the update step. Concretely, we calculate
$\rho=\frac{||F(\mathbf{x}+\Delta)||^{2}-||F(\mathbf{x})||^{2}}{||\mathbf{J}\Delta+F(\mathbf{x})||^{2}-||F(\mathbf{x})||^{2}}$ (5)
and only keep the update if $\rho>1e{-5}$, in which case we reduce the regularization strength $\lambda_{reg}$ by half. Otherwise we revert the update and double $\lambda_{reg}$.
3.3 Memory Efficient PCG Implementation
The PCG algorithm obtains the solution to the least squares problem of Eq. 3 in multiple iterations. In practice, we run the algorithm for $n_{iters}{=}8$ iterations and implement it with custom CUDA kernels. We summarize it in Algorithm 1.
1
Input : Gaussians and cameras $\mathcal{G}$, Residuals $\mathbf{F}$
Output : Update Direction $\mathbf{\Delta}$
2
[…]
12for $i=0$ to $n_{iters}$ do
13
14      $\mathbf{u}_{i}=\textnormal{{{\color[rgb]{0,0,1}applyJ}}}(\textnormal{{{\color[rgb]{0,0,1}sortX}}}(\mathbf{p}_{i}),\mathcal{G},\mathcal{C})$ // $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$
15      $\mathbf{g}_{i}=\textnormal{{{\color[rgb]{0,0,1}applyJT}}}(\mathbf{u}_{i},\mathcal{G},\mathcal{C})$ // $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$
16      $\alpha_{i}=\frac{\mathbf{r}_{i}^{T}\mathbf{z}_{i}}{\mathbf{p}_{i}^{T}\mathbf{g}_{i}}$
17      $\mathbf{x}_{i+1}{=}\mathbf{x}_{i}{+}\alpha_{i}\mathbf{p}_{i}$
18      $\mathbf{r}_{i+1}{=}\mathbf{r}_{i}{-}\alpha_{i}g_{i}$
19      $\mathbf{z}_{i+1}{=}\mathbf{M}^{-1}\mathbf{r}_{i+1}$
20      $\beta_{i}=\frac{\mathbf{r}^{T}_{i+1}\mathbf{z}_{i+1}}{\mathbf{r}^{T}_{i}\mathbf{z}_{i}}$
21      $\mathbf{p}_{i+1}=\mathbf{z}_{i+1}+\beta_{i}\mathbf{p}_{i}$
22 end for
23
return $\mathbf{x}_{i+1}$
Algorithm 1 We run the PCG algorithm with custom CUDA kernels (blue) in every LM iteration.
Most of the work in every PCG iteration is consumed by calculating the matrix-vector product $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{J}\mathbf{p}_{i}$. We implement this product in two stages by first calculating $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ and then $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$. Calculating the non-zero values of $\mathbf{J}$ requires backpropagating from every residual through the $\alpha$-blending and splat projection steps back to all Gaussian parameters. The original tile-based rasterizer of 3DGS [20] implements the calculation of $\mathbf{J}^{T}\mathbf{x}$ with a per-pixel parallelization. That is, every thread handles one ray, stepping backwards along all splats that this ray hit. We found that this parallelization is too slow for an efficient PCG implementation. The reason is that we have to repeat this calculation multiple times: per PCG iteration we do it once for $\mathbf{u}_{i}$ and once for $\mathbf{g}_{i}$. As a consequence, we re-calculate the same intermediate $\alpha$-blending states, since the loop over splats is evaluated multiple times.
Figure 3: Parallelization Strategy And Caching Scheme. We implement the PCG algorithm with efficient CUDA kernels, that use a gradient cache to calculate Jacobian-vector products. Before PCG starts, we create the gradient cache following the per-pixel parallelization of 3DGS [20]. Afterwards, we sort the cache by Gaussians to ensure coalesced read accesses. The cache decouples splats along rays, which allows us to parallelize per-pixel-per-splat during PCG.
Our key idea is to change the parallelization from per-pixel to per-pixel-per-splat. We summarize this pattern in Fig. 3. We describe the backward pass from a residual $r$ to a Gaussian parameter $x_{i}$ as:
$\frac{\partial r}{\partial x_{i}}=\frac{\partial r}{\partial p}\frac{\partial p}{\partial s}\frac{\partial s}{\partial x_{i}}$ (6)
where $\frac{\partial r}{\partial p}$ is the gradient from the residual to the pixel, $\frac{\partial p}{\partial s}$ from the pixel to the projected splat, and $\frac{\partial s}{\partial x_{i}}$ from the splat to the Gaussian parameter. Calculating $\frac{\partial p}{\partial s}$, requires access to the accumulated transmittance and color in the $\alpha$-blending process for that pixel until splat $s$. Instead of looping over all splats along a ray and recalculating $\frac{\partial p}{\partial s}$ every time, we cache this gradient once (Fig. 3 left). Every time we need to calculate $\mathbf{u}_{i}$ or $\mathbf{g}_{i}$ in the PCG algorithm, we then read the corresponding gradients from the cache (Fig. 3 right). This allows us to directly parallelize over all splats in all pixels, which drastically accelerates the runtime since we no longer have to loop over rays.
Storing these gradients in a cache consumes additional GPU memory and is largely controlled by how many images (rays) we process in each PCG iteration and how many splats contribute to the final color along each ray. We propose an efficient subsampling scheme in Sec. 3.3.2 to limit the memory to the available budget.
3.3.1 CUDA Kernel Design
We describe the parallelization pattern and output of every CUDA kernel that we use to implement Algorithm 1. We refer to the supplemental material for more details.
buildCache We use the per-pixel parallelization to calculate all gradients $\frac{\partial p}{\partial s}$. For coalesced write accesses, we store the cache values sorted by pixels (Fig. 3 left). Additionally, we calculate $\mathbf{b}=-\mathbf{J}^{T}\mathbf{F}$ following the implementation of the differentiable rasterizer in 3DGS [20], by splitting this calculation into three smaller kernels, where we write out the cache completely in the first kernel.
sortCacheByGaussians All remaining kernels require the cache to be sorted over Gaussians to ensure coalesced read accesses. We sort the cache by Gaussians and calculate a prefix sum over the number of entries per Gaussian. Subsequent kernels use this to index into the cache.
diagJTJ We use the Jacobi preconditioner $\mathbf{M}^{-1}{=}1/\text{diag}(\mathbf{J}^{T}\mathbf{J})$ when running PCG (Algorithm 1). This kernel computes the required entries by utilizing the per-pixel-per-splat parallelization and the cache. Every thread directly squares its calculated gradients and atomically adds them to the output vector.
applyJ We compute $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache. Additionally, we resort the input vector $\mathbf{p}_{i}$ for a coalesced memory access using the sortX kernel. Every thread sums up the calculated gradients and atomically adds them to the output vector.
applyJT We compute $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache. Similar to buildCache, we split the computation in three kernels.
3.3.2 Image Subsampling Scheme
Our caching data structure consumes additional GPU memory. For high resolution images in a dense reconstruction setup, the number of rays and thus the cache size can grow too large. To this end, we split the images into batches and solve the normal equations independently, following  Eq. 3. This allows us to store the cache only for one batch at a time. Concretely, for $n$ batches, we obtain $n$ update vectors and combine them in a weighted mean:
$\Delta=\sum_{i=1}^{n}\frac{\mathbf{M}_{i}\Delta_{i}}{\sum_{k=1}^{n}\mathbf{M}_{k}}$ (7)
where we use the inverse of the PCG preconditioner $M_{i}{=}\text{diag}(\mathbf{J}_{i}^{T}\mathbf{J}_{i})$ as the weights. We refer to the supplement for a derivation of the weights. These weights balance the importance of update vectors across batches based on how much each Gaussian parameter contributed to the final colors in the respective images. This subsampling scheme allows us to control the cache size relative to the number of images in a batch. In practice, we can choose batch sizes between 25 and 70 images and use up to $n{=}4$ batches per LM iteration. We either select the images at random or, if the scene was captured along a smooth trajectory, in a strided fashion to maximize the scene coverage in every batch.
3.4 3DGS Optimization in Two Stages
Figure 4: Comparison of initialization iterations. In our first stage, we initialize the Gaussians with gradient descent for $K$ iterations, before finetuning with our LM optimizer. After 6000 or 8000 iterations, our method converges faster than the baseline. With less iterations, pure LM is slower, which highlights the importance of our two stage approach. Results reported on the “garden” scene from MipNeRF360 [28] without densification.
We utilize our LM implementation in the second stage of 3DGS optimization (see Fig. 2). Before that, we use the ADAM optimizer to obtain an initialization of the Gaussian parameters. It is also possible to use the LM optimizer from the beginning, however this does not bring any additional speed-up (see Fig. 4). In the beginning of optimization, gradient descent makes rapid progress by optimizing the Gaussians from a single image per iteration. In contrast, we sample many images in every LM iteration, which makes every iteration more time-consuming. This additional compute overhead is especially helpful to converge to optimal Gaussian parameters quicker (see Fig. 1 left).
Splitting the method in two stages also allows us to complete the densification of the Gaussians before employing the LM optimizer, which simplifies the implementation.
4 Results
Baselines We compare our LM optimizer against ADAM in multiple reference implementations of 3DGS. This shows, that our method is compatible with other runtime improvements. In other words, we can swap out the optimizer and retain everything else. Concretely, we compare against the original 3DGS implementation [20], its reimplementation “gsplat” [41], and DISTWAR [11]. Additionally, we compare against Taming-3DGS [27] by utilizing their released source code, that contains the rasterizer improvements and uses the 3DGS [20] densification heuristics (referred to as “Taming-3DGS” in the following). We run all baselines for 30K iterations with their default hyperparameters.
Datasets and Metrics We benchmark our runtime improvements on three established datasets: Tanks&Temples [24], Deep Blending [18], and MipNeRF360 [4]. These datasets contain in total 13 scenes that cover bounded indoor and unbounded outdoor environments. We fit all scenes for every method on the same NVIDIA A100 GPU using the train/test split as proposed in the original 3DGS [20] publication. To measure the quality of the reconstruction, we report peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity (LPIPS) [44] averaged over all test images. Additionally, we report the optimization runtime and the maximum amount of consumed GPU memory.
Implementation Details For our main results, we run the first stage for 20K iterations with the default hyperparameters of the respective baseline. The densification is completed after 15K iterations. Afterwards, we only have to run 5 LM iterations with 8 PCG iterations each to converge on all scenes. This showcases the efficiency of our optimizer. Since the image resolutions are different for every dataset, we select the batch-size and number of batches such that the consumed memory for caching is similar. We select 25 images in 4 batches for MipNeRF360 [4], 25 images in 3 batches for Deep Blending [18], and 70 images in 3 batches for Tanks&Temples [24]. We constrain the value range of $\lambda_{reg}$ for stable updates. We define it in $[1e{-}4,1e{4}]$ for Deep Blending [18] and Tanks&Temples [24] and in the interval $[1e{-}4,1e{-}2]$ for MipNeRF360 [4].
4.1 Comparison to Baselines
Method MipNeRF-360 [4] Tanks&Temples [24] Deep Blending [18]
[…]
3DGS [20] 0.813 27.40 0.218 1271 0.844 23.68 0.178 736 0.900 29.51 0.247 1222
+ Ours 0.813 27.39 0.221 972 0.845 23.73 0.182 663 0.903 29.72 0.247 951
DISTWAR [11] 0.813 27.42 0.217 966 0.844 23.67 0.178 601 0.899 29.47 0.247 841
+ Ours 0.814 27.42 0.221 764 0.844 23.67 0.183 537 0.902 29.60 0.248 672
gsplat [41] 0.814 27.42 0.217 1064 0.846 23.50 0.179 646 0.904 29.52 0.247 919
+ Ours 0.814 27.42 0.221 818 0.844 23.68 0.183 414 0.902 29.58 0.249 716
Taming-3DGS [27] 0.810 27.44 0.224 684 0.847 23.79 0.176 463 0.902 29.71 0.242 590
+ Ours 0.814 27.41 0.221 589 0.844 23.72 0.183 393 0.902 29.72 0.249 540
Table 1: Quantitative comparison of our method and baselines. By adding our method to baselines, we accelerate the optimization time by 30% on average while achieving the same quality. We can combine our method with others, that improve runtime along different axes. This demonstrates that our method offers an orthogonal improvement, i.e., the LM optimizer can be plugged into many existing methods.
Refer to caption Refer to caption Refer to caption
3DGS [20] after 814s 3DGS + Ours after 794s Ground-Truth Images
Refer to caption Refer to caption Refer to caption
Taming-3DGS [27] after 456s Taming-3DGS + Ours after 454s Ground-Truth Images
Refer to caption Refer to caption Refer to caption
gsplat [41] after 453s gsplat + Ours after 447s Ground-Truth Images
Refer to caption Refer to caption Refer to caption
DISTWAR [11] after 978s DISTWAR + Ours after 971s Ground-Truth Images
Figure 5: Qualitative comparison of our method and baselines. We compare rendered test images after similar optimization time. All baselines converge faster when using our LM optimizer, which shows in images with fewer artifacts and more accurate brightness / contrast.
We report our main quantitative results in Tab. 1. Our LM optimizer can be added to all baseline implementations and accelerates the optimization runtime by 30% on average. The reconstructions show similar quality across all metrics and datasets, highlighting that our method arrives at similar local minima, just faster. We also provide a per-scene breakdown of these results in the supplemental material. On average our method consumes 53 GB of GPU memory on all datasets. In contrast, the baselines do not use an extra cache and only require between 6-11 GB of memory. This showcases the runtime-memory tradeoff of our approach.
We visualize sample images from the test set in Fig. 5 for both indoor and outdoor scenarios. After the same amount of optimization runtime, our method is already converged whereas the baselines still need to run longer. As a result, the baselines still contain suboptimal Gaussians, which results in visible artifacts in rendered images. In comparison, our rendered images more closely resemble the ground truth with more accurate brightness / contrast and texture details.
4.2 Ablations
Is the L1/SSIM objective important?
[…]
3DGS [20] (L1, SSIM) 0.862 27.23 0.108 1573
3DGS + Ours (L1, SSIM) 0.863 27.29 0.110 1175
3DGS [20] (L2) 0.854 27.31 0.117 1528
3DGS + Ours (L2) 0.857 27.48 0.114 1131
Table 2: Ablation of objective. We compare using the L1/SSIM losses against the L2 loss. For both, 3DGS [20] optimized with ADAM and combined with ours, we achieve better results with the L1/SSIM objective. In both cases, our method accelerates the convergence. We report results on the garden scene from the MipNeRF360 [4] dataset starting from the same initialization.
We utilize the same loss functions in our LM optimizer as in the original 3DGS implementation, namely the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses (see Eq. 1). Since LM energy terms are defined as a sum of squares, we adopt the square root formulation of these loss functions to arrive at an identical objective (see Eq. 2). We compare this choice against fitting the Gaussians with only an $\mathcal{L}_{2}$ loss, that does not require taking a square root. Concretely, we compare the achieved quality and runtime of LM against ADAM for both the $\mathcal{L}_{2}$ loss and the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses. As can be seen in Tab. 2, we achieve faster convergence and similar quality in both cases. However, the achieved quality is inferior for both LM and ADAM when only using the $\mathcal{L}_{2}$ loss. This highlights the importance of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ loss functions and why we adopt them in our method as well.
How many images per batch are necessary?
[…]
3DGS + Ours 5 100 0.969 33.77 0.030 242
3DGS + Ours 5 80 0.969 33.73 0.031 233
3DGS + Ours 5 60 0.968 33.69 0.031 223
3DGS + Ours 5 40 0.967 33.51 0.032 212
3DGS [20] 5 100 0.967 33.48 0.033 164
3DGS [20] 50 100 0.968 33.57 0.032 218
3DGS [20] 100 100 0.969 33.65 0.031 270
Table 3: Ablation of batch-size. We vary the number of iterations (#I) and number of images in a batch (#B). Reducing #B for our method reduces runtime and consumed memory, while only slightly impacting quality. This demonstrates that image subsampling Sec. 3.3.2 is compatible with LM in our task. Increasing #I in 3DGS [20] with ADAM is slower and achieves worse quality than Ours using LM. We report average results on the NeRF-Synthetic [28] dataset starting from the same initialization.
The key hyperparameters in our model are the number of images in a batch and how many batches to choose for every LM iteration (Sec. 3.3.2). This directly controls how much GPU memory our optimizer consumes and how much time we need per iteration. We compare different number of images in Tab. 3 (top) on the NeRF-Synthetic [28] dataset. We choose this dataset because it is possible to construct the cache for all 100 training images in 33 GB of GPU memory and optimize the scene with one batch of all images. We compare this against our proposed subsampling scheme for different number of images in a single batch. Decreasing the number of images in a batch results in only slightly worse quality, but also yields faster convergence and reduces GPU memory consumption linearly down to 15GB for 40 images. This demonstrates that subsampling images does not negatively impact the convergence of the LM optimizer in our task. This motivates our usage of subsampling on the real-world datasets in our main results.
Are we better than full-batch ADAM? One reason why LM requires drastically fewer iterations to converge than ADAM is the larger batch size (number of training images per iteration). Concretely, we require only 5 additional LM iterations after the initialization, whereas ADAM runs for another 10K iterations to converge. We can similarly increase the batch size in ADAM to reduce the number of iterations. However, as can be seen in Tab. 3 (bottom), the achieved quality is worse for the same batch size and number of iterations. When running for more iterations, ADAM eventually converges to similar quality, but needs more time. This highlights the efficiency of our optimizer: since we solve the normal equations in Eq. 2, one LM iteration makes a higher quality update step than ADAM which only uses the gradient direction.
4.3 Runtime Analysis
Figure 6: Runtime Analysis. One iteration of our LM optimizer is dominated by solving PCG and building the cache. Measured on the “garden” scene from Mip-NeRF360 [4] after densification.
We analyze the runtime of our LM optimizer across multiple iterations in Fig. 6. The runtime is dominated by solving Eq. 3 with PCG and building the cache (Sec. 3.3). Sorting the cache, rendering all selected images, and the line search (Eq. 4) are comparably faster. When running PCG, we execute the applyJ and applyJT kernels up to 8 times using the per-pixel-per-splat parallelization pattern (Algorithm 1). In contrast, we execute the buildCache kernel once and parallelize per-pixel, which is only marginally faster. This demonstrates the speed advantage obtained by using our proposed caching structure. We also provide a detailed profiling analysis of the kernels in the supplement.
4.4 Limitations
By replacing ADAM with our LM scheme, we accelerate the convergence speed by 30% on average for all datasets and baselines. However, some drawbacks remain.
First, our approach has a higher GPU memory footprint than baselines, due to our gradient cache (Sec. 3.3). Depending on the number of images and their resolution, this can make it hard to run our method on smaller GPUs. Following Mallick et al. [27], one could reduce the cache size by storing the gradients $\frac{\partial p}{\partial s}$ only for every 32nd splat along a ray and re-doing the $\alpha$-blending in these local windows.
Second, our two-stage approach relies on ADAM to first complete the densification. The original 3DGS [20] densifies Gaussians up to 140 times, which is not easily transferable to the granularity of only 5 LM iterations. Instead, one could explore recent improvements in densification and integrate them into our method [22, 5].
5 Conclusion
We have presented 3DGS-LM, a method that accelerates the reconstruction of 3D Gaussian-Splatting [20] by replacing the ADAM optimizer with a tailored Levenberg-Marquardt (LM) (Sec. 3.2). We show that with our data parallization scheme we can efficiently solve the normal equations with PCG in custom CUDA kernels (Sec. 3.3). Employed in a two-stage approach (Sec. 3.4), this leads to a 30% runtime acceleration compared to baselines. We further demonstrate that our approach is agnostic to other methods  [11, 41, 27], which further improves the optimization runtime. Overall, we believe that the ability of faster 3DGS reconstructions with our method will open up further research avenues and make existing 3DGS more practical across a wide range of real-world applications.
6 Acknowledgements
This project was funded by a Meta sponsored research agreement. In addition, the project was supported by the ERC Starting Grant Scan2CAD (804724) as well as the German Research Foundation (DFG) Research Unit “Learning and Simulation in Visual Computing”. We thank Justin Johnson for the helpful discussions in an earlier project with a similar direction and Peter Kocsis for the helpul discussions about image subsampling. We also thank Angela Dai for the video voice-over.
Supplementary Material
Appendix A More Details about CUDA Kernel Design
We introduce the necessary CUDA kernels to calculate the PCG algorithm in Sec. 3.3.1. In this section, we provide additional implementation details.
The necessary computations for the buildCache and applyJT steps in the PCG algorithm (see Algorithm 1) are split across three kernels. This follows the original design of the 3DGS differentiable rasterizer [20]. In both cases, we need to calculate the Jacobian-vector product of the form $\mathbf{g}=\mathbf{J}^{T}\mathbf{u}$ where $\mathbf{J}\in\mathbb{R}^{NxM}$ is the Jacobian matrix of $N$ residuals and $M$ Gaussian parameters and $\mathbf{u}\in\mathbb{R}^{N}$ is an input vector. The $k$-th element in the output vector is calculated as
$\mathbf{g}_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}\mathbf{u}_{i}=\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$ (8)
where $\mathbf{r}_{i}$ is the $i$-th residual and $\mathbf{x}_{k}$ is the $k$-th Gaussian parameter. Following the chain-rule, it is possible to split up the gradient $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}=\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}$. We can use this to split the computation across three smaller kernels, where only the first needs to calculate the sum over all residuals: $\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$. The other kernels then only calculate the remaining steps by parallelizing over Gaussians. Since the sum needs to be implemented atomically, i.e., multiple threads write to the same output position, this is the main bottleneck for the kernel implementation. By splitting the computation in three smaller parts, we only need to do the costly summation in the first kernel and only to intermediate attributes. Concretely $\mathbf{y}_{k}$ are the 2D mean, color, and opacity attributes of the $k$-th projected Gaussian. Since both the buildCache and applyJT kernels use the gradient cache and our proposed per-pixel-per-splat parallelization pattern, we can implement the sum by first doing a warp reduce and then only issuing one atomicAdd statement per warp.
In contrast, the applyJ and diagJTJ computations cannot be split up into smaller kernels. Concretely, the applyJ kernel calculates $\mathbf{u}=\mathbf{J}\mathbf{p}$ with $\mathbf{p}\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as
$\mathbf{u}_{k}=\sum_{i=0}^{M}\frac{\partial\mathbf{r}_{k}}{\partial\mathbf{x}_{i}}\mathbf{p}_{i}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{k}}{\partial\mathbf{y}_{i}}\frac{\partial\mathbf{y}_{i}}{\partial\mathbf{x}_{i}}\mathbf{p}_{i}$ (9)
Similarly, the diagJTJ kernel calculates $\mathbf{M}=\text{diag}(\mathbf{J}^{T}\mathbf{J})\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as
$\mathbf{g}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}})^{2}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}})^{2}$ (10)
In both cases it is not possible to move part of the gradients outside of the sum. As a consequence, both the applyJ and diagJTJ are implemented as one kernel, where each thread directly calculates the complete backward path to all Gaussian attributes. This slightly increases the number of required registers and the runtime compared to the applyJT kernel (see Tab. 4).
The applyJ kernel requires the input vector $\mathbf{p}$ to be sorted per Gaussian to make reading from it coalesced. That is: $\mathbf{p}=[x_{1}^{a},...,x_{1}^{z},...,x_{M}^{a},...,x_{M}^{z}]^{T}$, where $x_{k}^{a}$ is the value corresponding to the $a$-th parameter of the $k$-th Gaussian. In total, each Gaussian consists of 59 parameters: 11 for position, rotation, scaling, and opacity and 48 for all Spherical Harmonics coefficients of degree 3. In contrast, all other kernels require the input vector to be sorted per attribute to make reading from it coalesced. That is: $\mathbf{q}=[x_{1}^{a},...,x_{M}^{a},...,x_{1}^{z},...,x_{M}^{z}]^{T}$. We use the structure of $\mathbf{q}$ for all other vector-vector calculations in Algorithm 1 as well. Whenever we call the applyJ kernel, we thus first call the sortX kernel that restructures $\mathbf{q}$ to the layout of $\mathbf{p}$.
Appendix B Derivation of Weights for Subsampling
We sample batches of fewer images to decrease the size of the gradient cache (see Sec. 3.3.2). To combine the update vectors from multiple batches, we calculate the weighted mean, as detailed in Eq. 7. This weighted mean approximates the true solution without any image subsampling. That is, to obtain the update vector, we have to solve the normal equations, as detailed in Eq. 3. When subsampling images, we split the number of total residuals $M$ into smaller chunks. Let’s consider the case of two chunks, labeled as $1$, $2$. The normal equations (without subsampling) can then be written as:
$\begin{bmatrix}\mathbf{J}^{T}_{1}&\mathbf{J}^{T}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{J}_{1}\\ \mathbf{J}_{2}\end{bmatrix}\Delta=\begin{bmatrix}\mathbf{J}^{T}_{1}&\mathbf{J}^{T}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{F}_{1}(\mathbf{x})\\ \mathbf{F}_{2}(\mathbf{x})\end{bmatrix}$ (11)
where we drop the additional LM regularization term for clarity and divide the Jacobian and residual vector into separate matrices according to the chunks. The solution to the normal equations is obtained by:
$\Delta=(\mathbf{J}^{T}_{1}\mathbf{J}_{1}+\mathbf{J}^{T}_{2}\mathbf{J}_{2})^{-1}(\mathbf{J}^{T}_{1}\mathbf{F}_{1}(\mathbf{x})+\mathbf{J}^{T}_{2}\mathbf{F}_{2}(\mathbf{x}))$ (12)
In contrast, when we subsample images, we obtain two separate solutions as:
$\Delta_{1}=(\mathbf{J}^{T}_{1}\mathbf{J}_{1})^{-1}\mathbf{J}^{T}_{1}\mathbf{F}_{1}(\mathbf{x})$ (13)
$\Delta_{2}=(\mathbf{J}^{T}_{2}\mathbf{J}_{2})^{-1}\mathbf{J}^{T}_{2}\mathbf{F}_{2}(\mathbf{x})$ (14)
We can rewrite Eq. 12 as a weighted mean of $\Delta_{1}$, $\Delta_{2}$:
$\Delta=K^{-1}(\mathbf{J}^{T}_{1}\mathbf{J}_{1})(\mathbf{J}^{T}_{1}\mathbf{J}_{1})^{-1}(\mathbf{J}^{T}_{1}\mathbf{F}_{1}(\mathbf{x}))$ (15)
$+K^{-1}(\mathbf{J}^{T}_{2}\mathbf{J}_{2})(\mathbf{J}^{T}_{2}\mathbf{J}_{2})^{-1}(\mathbf{J}^{T}_{2}\mathbf{F}_{2}(\mathbf{x}))$ (16)
$=w_{1}\Delta_{1}+w_{2}\Delta_{2}$ (17)
where $K=(\mathbf{J}^{T}_{1}\mathbf{J}_{1}+\mathbf{J}^{T}_{2}\mathbf{J}_{2})$, $w_{1}=K^{-1}(\mathbf{J}^{T}_{1}\mathbf{J}_{1})$, $w_{2}=K^{-1}(\mathbf{J}^{T}_{2}\mathbf{J}_{2})$. Calculating these weights requires to materialize and invert $K$, which is too costly to fit in memory. To this end, we approximate the true weights $w_{1}$ and $w_{2}$ with $\tilde{w}_{1}=\text{diag}(w_{1})$ and $\tilde{w}_{2}=\text{diag}(w_{2})$. This directly leads to the weighted mean that we employ in Eq. 7.
Appendix C Detailed Analysis of Runtime
Kernel Duration (ms) $\downarrow$ Compute Throughput (%$\uparrow$ Memory Throughput (%$\uparrow$ Register Count $\downarrow$
buildCache_p1 31.32 78.56 78.56 64
buildCache_p2 0.53 17.43 87.94 58
buildCache_p3 4.12 4.54 73.45 74
sortCacheByGaussians 5.04 61.17 61.17 18
diagJTJ 41.60 71.13 71.13 90
sortX 4.45 15.15 60.30 36
applyJ 10.98 86.32 86.32 80
applyJT_p1 3.93 75.79 75.79 34
applyJT_p2 0.37 18.83 89.69 40
applyJT_p3 3.20 4.75 78.48 48
Table 4: Profiler Analysis of CUDA kernels. We provide results measured on a RTX3090 GPU for building/resorting the gradient cache and running one PCG iteration on the MipNerf360 [4] “garden” scene with a batch size of one image.
We provide additional analysis of the CUDA kernels by running the Nsight Compute [^1] https://developer.nvidia.com/nsight-compute profiler. We provide results in Tab. 4 measured on a RTX3090 GPU for building/resorting the gradient cache and running one PCG iteration on the MipNerf360 [4] “garden” scene with a batch size of one image. We add the suffixes _p1, _p2, _p3 to signal the three kernels that we use to implement the respective operation (see Appendix A).
Appendix D Results per Scene
Method Scene MipNeRF-360 [4]
[…]
3DGS [20] treehill 0.631 22.44 0.330 1130
+ Ours treehill 0.633 22.57 0.334 836
3DGS [20] counter 0.905 28.96 0.202 1178
+ Ours counter 0.904 28.89 0.206 927
3DGS [20] stump 0.769 26.56 0.217 1234
+ Ours stump 0.774 26.67 0.218 895
3DGS [20] bonsai 0.939 31.99 0.206 1034
+ Ours bonsai 0.938 31.84 0.208 794
3DGS [20] bicycle 0.764 25.20 0.212 1563
+ Ours bicycle 0.765 25.30 0.218 1141
3DGS [20] kitchen 0.925 31.37 0.128 1389
+ Ours kitchen 0.924 31.21 0.128 1156
3DGS [20] flowers 0.602 21.49 0.340 1132
+ Ours flowers 0.600 21.52 0.344 819
3DGS [20] room 0.917 31.36 0.221 1210
+ Ours room 0.916 31.10 0.224 1004
3DGS [20] garden 0.862 27.23 0.109 1573
+ Ours garden 0.863 27.30 0.110 1175
Method Scene Deep Blending [18]
[…]
3DGS [20] playroom 0.901 29.90 0.247 1085
+ Ours playroom 0.905 30.24 0.246 861
3DGS [20] drjohnson 0.898 29.12 0.246 1359
+ Ours drjohnson 0.901 29.23 0.248 1040
Method Scene Tanks & Temples [24]
[…]
3DGS [20] train 0.811 21.95 0.209 636
+ Ours train 0.811 22.07 0.214 579
3DGS [20] truck 0.877 25.40 0.148 837
+ Ours truck 0.876 25.36 0.151 747
Table 5: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
Method Scene MipNeRF-360 [4]
[…]
DISTWAR [11] treehill 0.633 22.47 0.327 898
+ Ours treehill 0.635 22.54 0.332 669
DISTWAR [11] counter 0.905 29.00 0.203 790
+ Ours counter 0.904 28.91 0.205 687
DISTWAR [11] stump 0.771 26.60 0.216 1017
+ Ours stump 0.773 26.70 0.217 760
DISTWAR [11] bonsai 0.939 32.13 0.206 677
+ Ours bonsai 0.938 31.92 0.208 578
DISTWAR [11] bicycle 0.763 25.19 0.212 1333
+ Ours bicycle 0.764 25.26 0.218 971
DISTWAR [11] kitchen 0.925 31.31 0.127 957
+ Ours kitchen 0.924 31.14 0.128 838
DISTWAR [11] flowers 0.602 21.45 0.340 884
+ Ours flowers 0.596 21.48 0.348 671
DISTWAR [11] room 0.916 31.41 0.221 803
+ Ours room 0.916 31.40 0.224 680
DISTWAR [11] garden 0.862 27.23 0.109 1338
+ Ours garden 0.861 27.32 0.112 1023
Method Scene Deep Blending [18]
[…]
DISTWAR [11] playroom 0.900 29.81 0.247 729
+ Ours playroom 0.905 30.24 0.246 586
DISTWAR [11] drjohnson 0.898 29.13 0.247 953
+ Ours drjohnson 0.901 29.13 0.249 758
Method Scene Tanks & Temples [24]
[…]
DISTWAR [11] train 0.812 22.05 0.209 504
+ Ours train 0.810 22.10 0.216 440
DISTWAR [11] truck 0.877 25.29 0.148 698
+ Ours truck 0.877 25.28 0.150 635
Table 6: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
Method Scene MipNeRF-360 [4]
[…]
gsplat [41] treehill 0.634 22.44 0.324 973
+ Ours treehill 0.635 22.54 0.332 701
gsplat [41] counter 0.908 28.99 0.201 903
+ Ours counter 0.904 28.91 0.205 762
gsplat [41] stump 0.769 26.53 0.218 1097
+ Ours stump 0.774 26.70 0.217 793
gsplat [41] bonsai 0.937 31.95 0.208 783
+ Ours bonsai 0.938 31.92 0.208 646
gsplat [41] bicycle 0.765 25.21 0.206 1398
+ Ours bicycle 0.765 25.26 0.218 988
gsplat [41] kitchen 0.926 31.17 0.128 1086
+ Ours kitchen 0.924 31.14 0.128 921
gsplat [41] flowers 0.600 21.53 0.338 965
+ Ours flowers 0.601 21.48 0.348 709
gsplat [41] room 0.920 31.48 0.219 913
+ Ours room 0.916 31.39 0.224 753
gsplat [41] garden 0.869 27.48 0.105 1462
+ Ours garden 0.861 27.32 0.112 1085
Method Scene Deep Blending [18]
[…]
gsplat [41] playroom 0.907 29.89 0.248 799
+ Ours playroom 0.904 30.90 0.247 626
gsplat [41] drjohnson 0.901 29.16 0.244 1040
+ Ours drjohnson 0.901 29.07 0.251 805
Method Scene Tanks & Temples [24]
[…]
gsplat [41] train 0.811 21.64 0.209 558
+ Ours train 0.809 22.09 0.216 381
gsplat [41] truck 0.880 25.35 0.149 735
+ Ours truck 0.877 25.28 0.150 447
Table 7: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
Method Scene MipNeRF-360 [4]
[…]
Taming-3DGS [27] treehill 0.630 22.61 0.340 638
+ Ours treehill 0.635 22.53 0.332 510
Taming-3DGS [27] counter 0.904 28.96 0.205 583
+ Ours counter 0.904 28.91 0.205 553
Taming-3DGS [27] stump 0.764 26.51 0.225 728
+ Ours stump 0.773 26.70 0.217 579
Taming-3DGS [27] bonsai 0.938 32.16 0.208 508
+ Ours bonsai 0.938 31.92 0.208 468
Taming-3DGS [27] bicycle 0.756 25.18 0.224 882
+ Ours bicycle 0.764 25.26 0.218 702
Taming-3DGS [27] kitchen 0.924 31.13 0.129 742
+ Ours kitchen 0.924 31.14 0.128 696
Taming-3DGS [27] flowers 0.594 21.42 0.351 620
+ Ours flowers 0.598 21.48 0.348 507
Taming-3DGS [27] room 0.916 31.64 0.224 561
+ Ours room 0.916 31.39 0.224 529
Taming-3DGS [27] garden 0.861 27.36 0.113 895
+ Ours garden 0.861 27.32 0.112 748
Method Scene Deep Blending [18]
[…]
Taming-3DGS [27] playroom 0.903 29.99 0.245 502
+ Ours playroom 0.904 30.23 0.247 453
Taming-3DGS [27] drjohnson 0.902 29.43 0.240 678
+ Ours drjohnson 0.901 29.21 0.249 627
Method Scene Tanks & Temples [24]
[…]
Taming-3DGS [27] train 0.815 22.18 0.205 411
+ Ours train 0.809 22.06 0.217 349
Taming-3DGS [27] truck 0.879 25.40 0.146 514
+ Ours truck 0.879 25.36 0.150 436
Table 8: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
We provide a per-scene breakdown of our main quantitative results against all baselines on all datasets. The comparisons against 3DGS [20] are in Tab. 5. The comparisons against DISTWAR [11] are in Tab. 6. The comparisons against gsplat [41] are in Tab. 7. The comparisons against Taming-3DGS [27] are in Tab. 8.

### v2

@@
 To this end, we change the optimizer to LM that runs in conjunction with the 3DGS differentiable rasterizer.
-For efficient GPU parallization, we propose a caching data structure for intermediate gradients that allows us to efficiently calculate Jacobian-vector products in custom CUDA kernels.
+For efficient GPU parallelization, we propose a caching data structure for intermediate gradients that allows us to efficiently calculate Jacobian-vector products in custom CUDA kernels.
 In every LM iteration, we calculate update directions from multiple image subsets using these kernels and combine them in a weighted mean.
-Overall, our method is 30% faster than the original 3DGS while obtaining the same reconstruction quality.
-Our optimization is also agnostic to other methods that acclerate 3DGS, thus enabling even faster speedups compared to vanilla 3DGS.
+Overall, our method is 20% faster than the original 3DGS while obtaining the same reconstruction quality.
+Our optimization is also agnostic to other methods that accelerate 3DGS, thus enabling even faster speedups compared to vanilla 3DGS.
 Refer to caption Refer to caption
 Figure 1: Our method accelerates 3D Gaussian Splatting (3DGS) [23] reconstruction by replacing the ADAM optimizer with a tailored Levenberg-Marquardt.
-Left: starting from the same initialization, our method converges 30% faster on the Tanks&Temples “Train” scene.
+Left: starting from the same initialization, our method converges faster on the Tanks&Temples TRAIN scene.
 Right: after the same amount of time, our method produces higher quality renderings (e.g., better brightness and contrast).
@@
 Existing methods reduce this runtime by improving the optimization along different axes.
-First, 3DGS renders images with a tile-based, differentiable rasterizer that is implemented in CUDA.
-In every iteration, the Gaussians get updated with gradient descent by backpropagating a rendering loss through the rasterizer.
-By improving the speed of the forward- and backward-pass, recent methods accelerate the optimization [11, 27, 41, 14].
+First, methods accelerate the rendering speed of the tile-based, differentiable rasterizer or the backward-pass that is specifically tailored for optimization with gradient descent [12, 32, 48, 15].
+For example, Durvasula et al[12] employ warp reductions for a more efficient sum of rendering gradients, while Mallick et al[32] utilizes a splat-parallelization for backpropagation.
 Second, in 3DGS the number of Gaussians is gradually grown during optimization, which is known as densification.
-To facilitate this, 3DGS accumulates positional gradients over multiple iterations and uses these statistics to split and prune Gaussians.
 Recently, GS-MCMC [25], Taming-3DGS [32], Mini-Splatting [14], and Revising-3DGS [5] propose novel densification schemes that reduce the number of required Gaussians to represent the scene.
-This makes the optimization more stable and also faster, since fewer Gaussians must be optimized.
+This makes the optimization more stable and also faster, since fewer Gaussians must be optimized and rendered in every iteration.
 Despite these improvements, the optimization still takes significant resources, requiring thousands of gradient descent iterations to converge.
@@
 More specifically, we propose to replace the widely used ADAM [26] optimizer with a tailored Levenberg-Marquardt (LM) [34].
+LM is known to drastically reduce the number of iterations by approximating second-order updates through solving the normal equations (Tab.
+4).
 This allows us to accelerate 3DGS reconstruction (Fig.
-1 left) by over 30% on average.
-Concretely, we propose a highly-efficient GPU parallization scheme for the preconditioned conjugate gradient (PCG) algorithm within the inner LM loop in order to obtain the respective update directions.
+1 left) by over 20% on average.
+Concretely, we propose a highly efficient GPU parallelization scheme for the preconditioned conjugate gradient (PCG) algorithm within the inner LM loop in order to obtain the respective update directions.
 To this end, we extend the differentiable 3DGS rasterizer with custom CUDA kernels that compute Jacobian-vector products.
@@
 In order to scale caching to high-resolution image datasets, we calculate update directions from multiple image subsets and combine them in a weighted mean.
-Overall, this allows us to improve reconstruction time by 30% compared to state-of-the-art 3DGS baselines while achiving the same reconstruction quality (Fig.
+Overall, this allows us to improve reconstruction time by 20% compared to state-of-the-art 3DGS baselines while achieving the same reconstruction quality (Fig.
 1 right).
 To summarize, our contributions are:
-  • we propose a tailored 3DGS optimization based on Levenberg-Marquard that improves reconstruction time by 30% and which is agnostic to other 3DGS accleration methods.
-  • we propose a highly-efficient GPU parallization scheme for the PCG algorithm for 3DGS in custom CUDA kernels with a caching data structure to facilitate efficient Jacobian-vector products.
+  • we propose a tailored 3DGS optimization based on Levenberg-Marquardt that improves reconstruction time by 20% and which is agnostic to other 3DGS acceleration methods.
+  • we propose a highly-efficient GPU parallelization scheme for the PCG algorithm for 3DGS in custom CUDA kernels with a caching data structure to facilitate efficient Jacobian-vector products.
 2 Related Work
@@
 Novel-View-Synthesis is widely explored in recent years [19, 2, 42, 33, 23, 3, 35].
-Neural Radiance Fields (NeRF) [28] are particularly successful, because they achieve highly photo-realistic image synthesis results by optimizing MLP weights through volume rendering.
-To improve training time from days to minutes, NeRF was combined with explicit representations like voxel grids [15, 35], hash grids [30], or points [40].
+NeRF [33] achieves highly photo-realistic image synthesis results through differentiable volumetric rendering.
+It was combined with explicit representations to accelerate optimization runtime [16, 41, 35, 47, 7].
 3D Gaussian Splatting (3DGS) [23] extends this idea by representing the scene as a set of 3D Gaussians, that are rasterized into 2D splats and then $\alpha$-blended into pixel colors.
-The approach gained popularity, due to the real-time rendering capabilities of high quality images.
-Since its inception, 3DGS was improved along several axes (next to improving optimization runtime).
-Recent methods improve the quality of rendered images [43, 17, 26] and the efficiency during rendering [31, 34].
-Others obtain better surface reconstructions [19, 16], reduce the memory requirements of the Gaussians [32], and enable training and rendering of large-scale scenes [21, 46].
-We similarly adopt 3DGS as our scene representation, but focus on improving the per-scene optimization time.
+The approach gained popularity, due to the ability to render high quality images in real-time.
+Since its inception, 3DGS was improved along several axes.
+Recent methods improve the image quality by increasing or regularizing the capacity of primitives [50, 18, 31, 20, 22].
+Others increase rendering efficiency [36, 40], obtain better surface reconstructions [21, 17], reduce the memory requirements [37], and enable large-scale reconstruction [24, 53].
+We similarly adopt 3DGS as our scene representation and focus on improving the per-scene optimization runtime.
 2.2 Speed-Up Gaussian Splatting Optimization
 Obtaining a 3DGS scene reconstruction can be accelerated in several ways.
-One line of work reduces the number of Gaussians by changing the densification heuristics [22, 27, 13, 5, 26].
-Recent methods focusing on sparse-view reconstruction train a neural network as data-driven prior, that directly outputs Gaussians in a single forward pass [12, 25, 8, 7, 6].
-In contrast, we focus on the dense-view and per-scene optimization setting, i.e., we do not require a data prior.
-Other works improve the optimization runtime by improving the implementation of the underlying differentiable rasterizer [41, 11, 27, 14].
-We demonstrate that our method is compatible with these approaches, i.e., our optimizer can be plugged into these methods to even further accelerate the optimization.
-2.3 Gauss-Newton Optimization For 3D Reconstruction Tasks
+One line of work reduces the number of Gaussians by changing the densification heuristics [25, 32, 14, 5, 31, 30].
+Other methods focus on sparse-view reconstruction and train a neural network as data prior, that outputs Gaussians in a single forward pass [13, 29, 9, 8, 6, 46, 54].
+In contrast, we focus on the dense-view and per-scene optimization setting, i.e., we are not limited to sparse-view reconstruction.
+Most related are methods that improve the implementation of the underlying differentiable rasterizer.
+In [12, 48] the gradient descent backward pass is accelerated through warp-reductions, while [32] improves its parallelization pattern and [15] accelerates the rendering.
+In contrast, we completely replace the gradient descent optimization with LM through a novel and tailored GPU parallelization scheme.
+We demonstrate that we are compatible with those existing methods, i.e., we further reduce runtime by plugging our optimizer into their scene initializations.
+2.3 Optimizers For 3D Reconstruction Tasks
 NeRF and 3DGS are typically optimized with stochastic gradient descent (SGD) optimizers like ADAM [26] for thousands of iterations.
@@
 Motivated by this, we aim to accelerate 3DGS optimization by adopting the Levenberg-Marquardt algorithm as our optimizer.
-Rasmuson et al.
-[33] implemented the Gauss-Newton algorithm for reconstructing a NeRF based on voxel grids.
-Their technical approach is related to ours, but we implement it for the 3DGS representation in a different way.
-Concretely, we subsample images in every iteration and introduce a caching data structure.
-This allows us to achieve state-of-the-art rendering quality, while significantly accelerating the optimization in comparison to existing methods.
+Rasmuson et al[39] implemented the Gauss-Newton algorithm for reconstructing low-resolution NeRFs based on dense voxel grids.
+In contrast, we exploit the explicit Gaussian primitives of 3DGS to perform highly-efficient Jacobian-vector products in a data-parallel fashion.
+This allows us to achieve state-of-the-art rendering quality, while significantly accelerating the optimization in comparison to ADAM-based methods.
 3 Method
@@
 The view-dependent color is modeled by Spherical Harmonics coefficients of order 3.
-To render an image of the scene from a specific viewpoint, all visible Gaussians are first projected into 2D splats with a tile-based differentiable rasterizer.
-Afterwards, the splats are $\alpha$-blended per-pixel to obtain a final pixel color.
-To fit the Gaussians to image observations, a rendering loss $\mathcal{L}$ is optimized with the ADAM [23] optimizer w.r.t.
-all Gaussian parameters $\mathbf{x}$:
-$\mathcal{L}(\mathbf{x}){=}\frac{1}{N}\sum_{i=1}^{N}(\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i}){+}\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i})))$ (1)
-where $\lambda_{1}{=}0.2$, $\lambda_{2}{=}0.8$, $\hat{c}_{i}$ is the rendered color, and $c_{i}$ the ground-truth color for one color channel of one pixel.
-Typically, ADAM uses a batch size of 1, meaning a random image per iteration is sampled to perform an update step.
-The number of Gaussians are initialized from the SfM points and is gradually grown during the first half of the optimization, which is known as densification [20].
+To render an image of the scene from a given viewpoint, all Gaussians are first projected into 2D Gaussian splats with a tile-based differentiable rasterizer.
+Afterwards, they are $\alpha$-blended along a ray to obtain the pixel color $c$:
+$c=\sum_{i\in\mathcal{N}}c_{i}\alpha_{i}T_{i},\quad\text{with }T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ (1)
+where $c_{i}$ is the color of the $i$-th splat along the ray, $\alpha_{i}$ is given by evaluating the 2D Gaussian multiplied with its opacity, and $T_{i}$ is the transmittance.
+To fit all Gaussian parameters $\mathbf{x}\in\mathbb{R}^{M}$ to posed image observations, a rendering loss is minimized with the ADAM [26] optimizer:
+$\mathcal{L}(\mathbf{x}){=}\frac{1}{N}\sum_{i=1}^{N}(\lambda_{1}|c_{i}{-}C_{i}|{+}\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i})))$ (2)
+where $\lambda_{1}{=}0.8$, $\lambda_{2}{=}0.2$, and $C_{i}$ the ground-truth for one pixel.
+Typically, 3DGS uses a batch size of 1 by sampling a random image per update step.
+The Gaussians are initialized from the SfM points and their number is gradually grown during the first half of the optimization, which is known as densification [23].
 3.2 Levenberg-Marquardt Optimization For 3DGS
 We employ the LM algorithm for optimization of the Gaussians by reformulating the rendering loss as a sum of squares energy function:
-$E(\mathbf{x}){=}\sum_{i=1}^{N}(\sqrt{\lambda_{1}\mathcal{L}_{1}(\hat{c}_{i},c_{i})}^{2}{+}\sqrt{\lambda_{2}(1{-}\mathcal{L}_{SSIM}(\hat{c}_{i},c_{i}))}^{2})$ (2)
-where we have in total $N{=}6HWP$ residuals for $P$ images with $H{x}W$ pixels each and 3 color channels.
-We take the square root of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses, to convert Eq.
+$E(\mathbf{x}){=}\sum_{i=1}^{N}\sqrt{\lambda_{1}|c_{i}{-}C_{i}|}^{2}{+}\sqrt{\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i}))}^{2}$ (3)
+where we have two separate residuals $r^{\text{abs}}_{i}{=}\sqrt{\lambda_{1}|c_{i}{-}C_{i}|}$ and $r^{\text{SSIM}}_{i}{=}\sqrt{\lambda_{2}(1{-}\text{SSIM}(c_{i},C_{i}))}$ per color channel of each pixel.
+We take the square root of each loss term, to convert Eq.
 2 into the required form for the LM algorithm.
@@
 In practice, we select large enough subsets of all images to ensure reliable update steps (see Sec.
-3.3.2 for more details).
-3DGS uses the structural similarity index measure (SSIM) as loss function during optimization.
-In SSIM, the local neighborhood of every pixel gets convolved with Gaussian kernels to obtain the final per-pixel score [39].
-As a consequence, gradients from the loss flow from every pixel to all Gaussians in the local neighborhood of that pixel.
-In contrast, the $\mathcal{L}_{1}$ loss only provides gradients from every pixel to the Gaussians along the corresponding ray.
-For simplicity, we approximate the gradient flow in $\mathcal{L}_{SSIM}$ by backpropagating the per-pixel scores only to the center pixels (ignoring the contribution to other pixels in the local neighborhood).
-We implement it following the derivation of Zhao et al.
-[45].
-This approximation allows us to keep rays independent from each other when calculating the Jacobian, which is a desirable property for our CUDA kernel implementation (see Sec.
 3.3 for more details).
-Obtaining Update Directions
-In every iteration of our optimization we obtain the update direction $\Delta\in\mathbb{R}^{M}$ for all $M$ Gaussian parameters by solving the normal equations:
-$(\mathbf{J}^{T}\mathbf{J}+\lambda_{reg}\textit{diag}(\mathbf{J}^{T}\mathbf{J}))\Delta=\mathbf{J}^{T}\mathbf{F}(\mathbf{x})$ (3)
-where $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{N}$ is the residual vector corresponding to Eq. 2 and $\mathbf{J}\in\mathbb{R}^{NxM}$ the corresponding Jacobian matrix.
-In a typical dense capture setup, we optimize over millions of Gaussians and have hundreds of high-resolution images.
+Obtaining Update Directions In every iteration of our optimization we obtain the update direction $\Delta\in\mathbb{R}^{M}$ for all $M$ Gaussian parameters by solving the normal equations:
+$(\mathbf{J}^{T}\mathbf{J}+\lambda_{\text{reg}}\text{diag}(\mathbf{J}^{T}\mathbf{J}))\Delta=-\mathbf{J}^{T}\mathbf{F}(\mathbf{x})$ (4)
+where $\mathbf{F}(\mathbf{x}){=}[r^{\text{abs}}_{1},...,r^{\text{abs}}_{N},r^{\text{SSIM}}_{1},...,r^{\text{SSIM}}_{N}]{\in}\mathbb{R}^{2N}$ is the residual vector corresponding to Eq.
+3 and $\mathbf{J}{\in}\mathbb{R}^{2N{\times}M}$ the corresponding Jacobian matrix.
+In a typical dense capture setup, we optimize over millions of Gaussians and have hundreds of high-resolution images [27, 19, 4].
 Even though $\mathbf{J}$ is a sparse matrix (each row only contains non-zero values for the Gaussians that contribute to the color of that pixel), it is therefore not possible to materialize $\mathbf{J}$ in memory.
@@
 3.3 for more details.
-Apply Parameter Update
-After we obtained the solution $\Delta$, we run a line search to find the best scaling factor $\gamma$ for updating the Gaussian parameters:
-$\min_{\gamma}E(\mathbf{x}_{k}-\gamma\Delta)$ (4)
+Apply Parameter Update After we obtained the solution $\Delta$, we run a line search to find the best scaling factor $\gamma{\in}\mathbb{R}$ for updating the Gaussian parameters:
+$\min_{\gamma}E(\mathbf{x}_{k}+\gamma\Delta)$ (5)
 In practice, we run the line search on a 30% subset of all images, which is enough to get a reasonable estimate for $\gamma$, but requires fewer rendering passes.
-Afterwards, we update the Gaussian parameters as: $\mathbf{x}_{k+1}{=}\mathbf{x}_{k}-\gamma\Delta$.
-Similar to the implementation of LM in CERES [1], we adjust the regularization strength $\lambda_{reg}$ after every iteration based on the quality of the update step.
+Afterwards, we update the Gaussian parameters as: $\mathbf{x}_{k+1}{=}\mathbf{x}_{k}+\gamma\Delta$.
+Similar to the implementation of LM in CERES [1], we adjust the regularization strength $\lambda_{\text{reg}}{\in}\mathbb{R}$ after every iteration based on the quality of the update step.
 Concretely, we calculate
-$\rho=\frac{||F(\mathbf{x}+\Delta)||^{2}-||F(\mathbf{x})||^{2}}{||\mathbf{J}\Delta+F(\mathbf{x})||^{2}-||F(\mathbf{x})||^{2}}$ (5)
-and only keep the update if $\rho>1e{-5}$, in which case we reduce the regularization strength $\lambda_{reg}$ by half. Otherwise we revert the update and double $\lambda_{reg}$.
-3.3 Memory Efficient PCG Implementation
+[…]
+and only keep the update if $\rho{>}1\mathrm{e}{-5}$, in which case we reduce the regularization strength as $\lambda_{\text{reg}}{\mathrel{*}=}1{-}(2\rho-1)^{3}$.
+Otherwise, we revert the update and double $\lambda_{\text{reg}}$.
+3.3 Efficient Parallelization Scheme For PCG
 The PCG algorithm obtains the solution to the least squares problem of Eq.
 4 in multiple iterations.
-In practice, we run the algorithm for $n_{iters}{=}8$ iterations and implement it with custom CUDA kernels.
+We run the algorithm for up to $\text{n}_{\text{iters}}{=}8$ iterations and implement it with custom CUDA kernels.
 We summarize it in Algorithm 1.
 1
-Input : Gaussians and cameras $\mathcal{G}$, Residuals $\mathbf{F}$
-Output : Update Direction $\mathbf{\Delta}$
+Input : Gaussians and cameras $\mathcal{G}$, $\mathbf{F}$, $\lambda_{\text{reg}}$
+Output : Update direction $\Delta$
 2
@@
 […]
-12for $i=0$ to $n_{iters}$ do
+12for $i=0$ to $\text{n}_{\text{iters}}$ do
 13
@@
 15 $\mathbf{g}_{i}=\textnormal{{{\color[rgb]{0,0,1}applyJT}}}(\mathbf{u}_{i},\mathcal{G},\mathcal{C})$ // $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$
+16 $\mathbf{g}_{i}\mathrel{+}=\lambda_{\text{reg}}\mathbf{M}\mathbf{p}_{i}$
 17 $\alpha_{i}=\frac{\mathbf{r}_{i}^{T}\mathbf{z}_{i}}{\mathbf{p}_{i}^{T}\mathbf{g}_{i}}$
 18 $\mathbf{x}_{i+1}{=}\mathbf{x}_{i}{+}\alpha_{i}\mathbf{p}_{i}$
-18      $\mathbf{r}_{i+1}{=}\mathbf{r}_{i}{-}\alpha_{i}g_{i}$
+19 $\mathbf{r}_{i+1}{=}\mathbf{r}_{i}{-}\alpha_{i}\mathbf{g}_{i}$
 20 $\mathbf{z}_{i+1}{=}\mathbf{M}^{-1}\mathbf{r}_{i+1}$
@@
 22 $\mathbf{p}_{i+1}=\mathbf{z}_{i+1}+\beta_{i}\mathbf{p}_{i}$
+[…]
+24     break
+25   end if
+26
 27 end for
-23
+28
 return $\mathbf{x}_{i+1}$
@@
 Most of the work in every PCG iteration is consumed by calculating the matrix-vector product $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{J}\mathbf{p}_{i}$.
-We implement this product in two stages by first calculating $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ and then $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$.
-Calculating the non-zero values of $\mathbf{J}$ requires backpropagating from every residual through the $\alpha$-blending and splat projection steps back to all Gaussian parameters.
-The original tile-based rasterizer of 3DGS [20] implements the calculation of $\mathbf{J}^{T}\mathbf{x}$ with a per-pixel parallelization.
+We compute it by first calculating $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ and then $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$.
+Calculating the non-zero values of $\mathbf{J}$ requires backpropagating from the residuals through the $\alpha$-blending (Eq.
+1) and splat projection steps to the Gaussian parameters.
+The tile-based rasterizer of 3DGS [23] performs this calculation using a per-pixel parallelization.
 That is, every thread handles one ray, stepping backwards along all splats that this ray hit.
 We found that this parallelization is too slow for an efficient PCG implementation.
-The reason is that we have to repeat this calculation multiple times: per PCG iteration we do it once for $\mathbf{u}_{i}$ and once for $\mathbf{g}_{i}$.
-As a consequence, we re-calculate the same intermediate $\alpha$-blending states, since the loop over splats is evaluated multiple times.
+The reason is the repetition of the ray marching: per PCG iteration we do it once for $\mathbf{u}_{i}$ and once for $\mathbf{g}_{i}$.
+As a consequence, the same intermediate $\alpha$-blending states (i.e., $T_{s}$, $\frac{\partial c}{\partial\alpha_{s}}$, $\frac{\partial c}{\partial c_{s}}$ for every splat $s$ along the ray) are re-calculated multiple (up to 18) times.
+Cache-driven parallelization
 Figure 3: Parallelization Strategy And Caching Scheme.
 We implement the PCG algorithm with efficient CUDA kernels, that use a gradient cache to calculate Jacobian-vector products.
-Before PCG starts, we create the gradient cache following the per-pixel parallelization of 3DGS [20].
+Left: before PCG starts, we create the gradient cache following the per-pixel parallelization of 3DGS [23].
 Afterwards, we sort the cache by Gaussians to ensure coalesced read accesses.
-The cache decouples splats along rays, which allows us to parallelize per-pixel-per-splat during PCG.
-Our key idea is to change the parallelization from per-pixel to per-pixel-per-splat.
-We summarize this pattern in Fig.
-3.
-We describe the backward pass from a residual $r$ to a Gaussian parameter $x_{i}$ as:
-$\frac{\partial r}{\partial x_{i}}=\frac{\partial r}{\partial p}\frac{\partial p}{\partial s}\frac{\partial s}{\partial x_{i}}$ (6)
-where $\frac{\partial r}{\partial p}$ is the gradient from the residual to the pixel, $\frac{\partial p}{\partial s}$ from the pixel to the projected splat, and $\frac{\partial s}{\partial x_{i}}$ from the splat to the Gaussian parameter.
-Calculating $\frac{\partial p}{\partial s}$, requires access to the accumulated transmittance and color in the $\alpha$-blending process for that pixel until splat $s$.
-Instead of looping over all splats along a ray and recalculating $\frac{\partial p}{\partial s}$ every time, we cache this gradient once (Fig.
+Right: the cache decouples splats along rays, which allows us to parallelize per-pixel-per-splat when computing $\mathbf{u}=\mathbf{J}\mathbf{p}$ and $\mathbf{g}=\mathbf{J}^{T}\mathbf{u}$ during PCG.
+We propose to change the parallelization to per-pixel-per-splat (summarized in Fig.
+3).
+That is, one thread handles all residuals of one ray for one splat.
+Each entry of $\mathbf{J}$ is the gradient from a residual $r$ (either of the L1 or SSIM terms) to a Gaussian parameter $x_{i}$.
+Conceptually, this can be computed in three stages:
+$\frac{\partial r}{\partial x_{i}}=\frac{\partial r}{\partial c}\frac{\partial c}{\partial s}\frac{\partial s}{\partial x_{i}}$ (7)
+where $\frac{\partial r}{\partial c}$ denotes the gradient from the residual to the rendered color, $\frac{\partial c}{\partial s}$ from the color to the projected splat, and $\frac{\partial s}{\partial x_{i}}$ from the splat to the Gaussian parameter.
+The first and last factors of Eq.
+7 can be computed independently for each residual and splat respectively, which allows for an efficient parallelization.
+Similarly, we can calculate $\frac{\partial c}{\partial s}$ independently, if we have access to $T_{s}$ and $\frac{\partial c}{\partial\alpha_{s}}$.
+Instead of looping over all splats along a ray multiple times, we cache these quantities once (Fig.
 3 left).
-Every time we need to calculate $\mathbf{u}_{i}$ or $\mathbf{g}_{i}$ in the PCG algorithm, we then read the corresponding gradients from the cache (Fig.
+When calculating $\mathbf{u}_{i}$ or $\mathbf{g}_{i}$, we then read these values from the cache (Fig.
 3 right).
-This allows us to directly parallelize over all splats in all pixels, which drastically accelerates the runtime since we no longer have to loop over rays.
-Storing these gradients in a cache consumes additional GPU memory and is largely controlled by how many images (rays) we process in each PCG iteration and how many splats contribute to the final color along each ray.
-We propose an efficient subsampling scheme in Sec.
-3.3.2 to limit the memory to the available budget.
-3.3.1 CUDA Kernel Design
-We describe the parallelization pattern and output of every CUDA kernel that we use to implement Algorithm 1. We refer to the supplemental material for more details.
-buildCache We use the per-pixel parallelization to calculate all gradients $\frac{\partial p}{\partial s}$.
-For coalesced write accesses, we store the cache values sorted by pixels (Fig.
+This allows us to parallelize over all splats in all pixels, which drastically accelerates the runtime.
+The cache size is controlled by how many images (rays) we process in each PCG iteration and how many splats contribute to the final color along each ray.
+We propose an efficient subsampling scheme that limits the cache size to the available budget.
+3DGS uses the structural similarity index measure (SSIM) as loss term during optimization (Eq.
+2).
+In SSIM, the local neighborhood of every pixel gets convolved with Gaussian kernels to obtain the final per-pixel score [45].
+We calculate $\frac{\partial r}{\partial c}$ for the SSIM residuals by backpropagating the per-pixel scores to the center pixels (ignoring the contribution to other pixels in the local neighborhood).
+This allows us to keep rays independent of each other thereby allowing for an efficient parallelization.
+We implement it following the derivation of Zhao et al[52].
+Mapping of PCG to CUDA kernels We cache all gradients $\frac{\partial c}{\partial s}$ using the buildCache operation.
+Following the implementation of the differentiable rasterizer in 3DGS [23], it uses the per-pixel parallelization and calculates the gradient update $\mathbf{b}{=}-\mathbf{J}^{T}\mathbf{F}$.
+For coalesced read and write accesses, we first store the cache sorted by pixels (Fig.
 3 left).
-Additionally, we calculate $\mathbf{b}=-\mathbf{J}^{T}\mathbf{F}$ following the implementation of the differentiable rasterizer in 3DGS [20], by splitting this calculation into three smaller kernels, where we write out the cache completely in the first kernel.
-sortCacheByGaussians All remaining kernels require the cache to be sorted over Gaussians to ensure coalesced read accesses.
-We sort the cache by Gaussians and calculate a prefix sum over the number of entries per Gaussian.
-Subsequent kernels use this to index into the cache.
-diagJTJ We use the Jacobi preconditioner $\mathbf{M}^{-1}{=}1/\text{diag}(\mathbf{J}^{T}\mathbf{J})$ when running PCG (Algorithm 1).
-This kernel computes the required entries by utilizing the per-pixel-per-splat parallelization and the cache.
-Every thread directly squares its calculated gradients and atomically adds them to the output vector.
-applyJ We compute $\mathbf{u}_{i}{=}\mathbf{J}\mathbf{p}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache.
-Additionally, we resort the input vector $\mathbf{p}_{i}$ for a coalesced memory access using the sortX kernel.
-Every thread sums up the calculated gradients and atomically adds them to the output vector.
-applyJT We compute $\mathbf{g}_{i}{=}\mathbf{J}^{T}\mathbf{u}_{i}$ by utilizing the per-pixel-per-splat parallelization and the cache. Similar to buildCache, we split the computation in three kernels.
-3.3.2 Image Subsampling Scheme
-Our caching data structure consumes additional GPU memory.
+Afterwards, we re-sort it by Gaussians using the sortCacheByGaussians kernel.
+We use the Jacobi preconditioner $\mathbf{M}^{-1}{=}1/\text{diag}(\mathbf{J}^{T}\mathbf{J})$ and calculate it once using the per-pixel-per-splat parallelization in the diagJTJ kernel.
+The inner PCG loop involves two kernels that are accelerated by our novel parallelization scheme.
+First, applyJ computes $\mathbf{u}{=}\mathbf{J}\mathbf{p}$, which we implement as a per-pixel sum aggregation.
+Afterwards, applyJT computes $\mathbf{g}{=}\mathbf{J}^{T}\mathbf{u}$.
+This per-Gaussian sum can be efficiently aggregated using warp reductions.
+We compute the remaining vector-vector terms of Algorithm 1 directly in PyTorch [38].
+We refer to the supplementary material for more details.
+Image Subsampling Scheme Our cache consumes additional GPU memory.
 For high resolution images in a dense reconstruction setup, the number of rays and thus the cache size can grow too large.
 To this end, we split the images into batches and solve the normal equations independently, following  Eq.
-3.
+4.
 This allows us to store the cache only for one batch at a time.
-Concretely, for $n$ batches, we obtain $n$ update vectors and combine them in a weighted mean:
-$\Delta=\sum_{i=1}^{n}\frac{\mathbf{M}_{i}\Delta_{i}}{\sum_{k=1}^{n}\mathbf{M}_{k}}$ (7)
-where we use the inverse of the PCG preconditioner $M_{i}{=}\text{diag}(\mathbf{J}_{i}^{T}\mathbf{J}_{i})$ as the weights.
-We refer to the supplement for a derivation of the weights.
-These weights balance the importance of update vectors across batches based on how much each Gaussian parameter contributed to the final colors in the respective images.
+Concretely, for $\text{n}_{\text{b}}$ batches, we obtain $\text{n}_{\text{b}}$ update vectors and combine them in a weighted mean:
+$\Delta=\sum_{i=1}^{\text{n}_{\text{b}}}\frac{\mathbf{M}_{i}\Delta_{i}}{\sum_{k=1}^{n}\mathbf{M}_{k}}$ (8)
+where we use the inverse of the PCG preconditioner $\mathbf{M}_{i}{=}\text{diag}(\mathbf{J}_{i}^{T}\mathbf{J}_{i})$ as the weights.
+We refer to the supplementary material for a derivation of the weights.
+These weights balance the importance of update vectors across batches based on how much each Gaussian parameter contributed to the rendered colors in the respective images.
 This subsampling scheme allows us to control the cache size relative to the number of images in a batch.
-In practice, we can choose batch sizes between 25 and 70 images and use up to $n{=}4$ batches per LM iteration.
-We either select the images at random or, if the scene was captured along a smooth trajectory, in a strided fashion to maximize the scene coverage in every batch.
+In practice, we choose batch sizes of 25-70 images and up to $\text{n}_{\text{b}}{=}4$ batches per LM iteration.
+We either select the images at random or, if the scene was captured along a smooth trajectory, in a strided fashion to maximize scene coverage in all batches.
 3.4 3DGS Optimization In Two Stages
+Our pipeline utilizes the LM optimizer in the second stage of 3DGS optimization (see Fig.
+2).
+Before that, we run the ADAM optimizer to obtain an initialization of the Gaussian parameters.
+We compare this against running our LM optimizer directly on the Gaussian initialization obtained from the SfM point cloud (following [23]).
+Fig.
+4 shows, that our LM converges faster for better initialized Gaussians and eventually beats pure ADAM.
+In contrast, running it directly on the SfM initialization is slower.
+This demonstrates that quasi second-order solvers like ours are well-known to be more sensitive to initialization.
+In other words, gradient descent makes rapid progress in the beginning, but needs more time to converge to final Gaussian parameters.
+The additional compute overhead of our LM optimization is especially helpful to converge more quickly.
+This motivates us to split the method in two stages.
+It also allows us to complete the densification of the Gaussians before employing the LM optimizer, which simplifies the implementation.
+4 Results
 Figure 4: Comparison of initialization iterations.
 In our first stage, we initialize the Gaussians with gradient descent for K iterations, before finetuning with our LM optimizer.
-After 6000 or 8000 iterations, our method converges faster than the baseline.
+After $\text{K}{=}6000$ or $\text{K}{=}8000$ iterations, our method converges faster than the baseline.
 With less iterations, pure LM is slower, which highlights the importance of our two stage approach.
-Results reported on the “garden” scene from MipNeRF360 [28] without densification.
-We utilize our LM implementation in the second stage of 3DGS optimization (see Fig.
-2).
-Before that, we use the ADAM optimizer to obtain an initialization of the Gaussian parameters.
-It is also possible to use the LM optimizer from the beginning, however this does not bring any additional speed-up (see Fig.
-4).
-In the beginning of optimization, gradient descent makes rapid progress by optimizing the Gaussians from a single image per iteration.
-In contrast, we sample many images in every LM iteration, which makes every iteration more time-consuming.
-This additional compute overhead is especially helpful to converge to optimal Gaussian parameters quicker (see Fig.
-1 left).
-Splitting the method in two stages also allows us to complete the densification of the Gaussians before employing the LM optimizer, which simplifies the implementation.
-4 Results
+Results reported on the GARDEN scene from MipNeRF360 [33] without densification.
 Baselines We compare our LM optimizer against ADAM in multiple reference implementations of 3DGS.
@@
 In other words, we can swap out the optimizer and retain everything else.
-Concretely, we compare against the original 3DGS implementation [20], its reimplementation “gsplat” [41], and DISTWAR [11].
-Additionally, we compare against Taming-3DGS [27] by utilizing their released source code, that contains the rasterizer improvements and uses the 3DGS [20] densification heuristics (referred to as “Taming-3DGS” in the following).
+Concretely, we compare against the original 3DGS [23], its reimplementation “gsplat” [48], and DISTWAR [12].
+Additionally, we compare against Taming-3DGS [32] by utilizing their “budgeted” approach as the fastest baseline in terms of runtime.
 We run all baselines for 30K iterations with their default hyperparameters.
-Datasets and Metrics We benchmark our runtime improvements on three established datasets: Tanks&Temples [24], Deep Blending [18], and MipNeRF360 [4].
+Datasets / Metrics We benchmark our runtime improvements on three established datasets: Tanks&Temples [27], Deep Blending [19], and MipNeRF360 [4].
 These datasets contain in total 13 scenes that cover bounded indoor and unbounded outdoor environments.
@@
 We select 25 images in 4 batches for MipNeRF360 [4], 25 images in 3 batches for Deep Blending [19], and 70 images in 3 batches for Tanks&Temples [27].
-We constrain the value range of $\lambda_{reg}$ for stable updates.
-We define it in $[1e{-}4,1e{4}]$ for Deep Blending [18] and Tanks&Temples [24] and in the interval $[1e{-}4,1e{-}2]$ for MipNeRF360 [4].
+We constrain the value range of $\lambda_{\text{reg}}$ for stable updates.
+We define it in $1\mathrm{e}{-4},1\mathrm{e}{4}]$ for Deep Blending [19] and Tanks&Temples [27] and in the interval $1\mathrm{e}{-4},1\mathrm{e}{-2}]$ for MipNeRF360 [4].
 4.1 Comparison To Baselines
@@
 + Ours 0.814 27.42 0.221 818 0.844 23.68 0.183 414 0.902 29.58 0.249 716
-Taming-3DGS [27] 0.810 27.44 0.224 684 0.847 23.79 0.176 463 0.902 29.71 0.242 590
-+ Ours 0.814 27.41 0.221 589 0.844 23.72 0.183 393 0.902 29.72 0.249 540
+Taming-3DGS [32] 0.793 27.14 0.260 566 0.833 23.76 0.209 366 0.900 29.84 0.274 447
++ Ours 0.791 27.13 0.260 453 0.832 23.72 0.209 310 0.901 29.91 0.275 347
 Table 1: Quantitative comparison of our method and baselines.
-By adding our method to baselines, we accelerate the optimization time by 30% on average while achieving the same quality.
+By adding our method to baselines, we accelerate the optimization time by 20% on average while achieving the same quality.
 We can combine our method with others, that improve runtime along different axes.
@@
 Refer to caption Refer to caption Refer to caption
-Taming-3DGS [27] after 456s Taming-3DGS + Ours after 454s Ground-Truth Images
+Taming-3DGS [32] after 328s Taming-3DGS + Ours after 324s Ground-Truth Images
 Refer to caption Refer to caption Refer to caption
@@
 1.
-Our LM optimizer can be added to all baseline implementations and accelerates the optimization runtime by 30% on average.
+Our LM optimizer can be added to all baseline implementations and accelerates the optimization runtime by 20% on average.
 The reconstructions show similar quality across all metrics and datasets, highlighting that our method arrives at similar local minima, just faster.
-We also provide a per-scene breakdown of these results in the supplemental material.
+We also provide a per-scene breakdown of these results in the supplementary material.
 On average our method consumes 53 GB of GPU memory on all datasets.
@@
 Is the L1/SSIM objective important?
+We utilize the same objective in our LM optimizer as in the original 3DGS implementation, namely the L1 and SSIM loss terms (Eq.
+2).
+Since LM energy terms are defined as a sum of squares, we adopt the square root formulation of these loss terms to arrive at an identical objective (Eq.
+3).
+We compare this choice against fitting the Gaussians with only an L2 loss, that does not require taking a square root.
+Concretely, we compare the achieved quality and runtime of LM against ADAM for both the L2 loss and the L1 and SSIM losses.
+As can be seen in Tab.
+2, we achieve faster convergence and similar quality in both cases.
+However, the achieved quality is inferior for both LM and ADAM when only using the L2 loss.
+This highlights the importance of the L1 and SSIM loss terms and why we adopt them in our method as well.
+We show in the supplementary material, that computing these loss terms instead of the simpler L2 residuals does not negatively impact the efficiency of our CUDA kernels.
+How many images per batch are necessary?
+The key hyperparameters in our model are the number of images in a batch and how many batches to choose for every LM iteration (Sec.
+3.3).
+This controls the runtime of one iteration and how much GPU memory our optimizer consumes.
+We compare different numbers of images in Tab.
+3 on the NeRF-Synthetic [33] dataset in a single batch per LM iteration, i.e., $\text{n}_{\text{b}}{=}1$.
+Using the full dataset (100 images) produces the best results.
+Decreasing the number of images in a batch results in only slightly worse quality, but also yields faster convergence and reduces GPU memory consumption linearly down to 15GB for 40 images.
+This demonstrates that subsampling images does not negatively impact the convergence of the LM optimizer in our task.
+Are we better than multi-view ADAM?
+Our method converges with fewer iterations than baselines.
+Concretely, we require only 5-10 additional LM iterations after the initialization, whereas ADAM runs for another 10K iterations.
+We increase the batch-size (number of images) for the baselines, such that the same number of multi-view constraints are observed for the respective update steps.
+However, as can be seen in Tab.
+4, the achieved quality is worse for ADAM after the same number of iterations.
+When running for more iterations, ADAM eventually converges to similar quality, but needs more time.
+This highlights the efficiency of our optimizer: since we solve the normal equations in Eq.
+3, one LM iteration makes a higher quality update step than ADAM which only uses the gradient direction.
+4.3 Runtime Analysis
+We analyze the runtime of our LM optimizer across multiple iterations in Fig.
+6.
+The runtime is dominated by solving Eq.
+4 with PCG and building the cache (Sec.
+3.3).
+Sorting the cache, rendering the selected images, and the line search (Eq.
+5) are comparatively faster.
+During PCG, we run the applyJ and applyJT kernels up to 9 times, parallelizing per-pixel-per-splat.
+In contrast, we run the buildCache kernel once, parallelizing per-pixel, which is only marginally faster.
+This shows the advantage of our proposed parallelization scheme: the same Jacobian-vector product runs much faster.
+We also provide a detailed profiling analysis of our kernels in the supplementary material.
+4.4 Limitations
 […]
-3DGS [20] (L1, SSIM) 0.862 27.23 0.108 1573
-3DGS + Ours (L1, SSIM) 0.863 27.29 0.110 1175
+3DGS [23] (L1/SSIM) 0.862 27.23 0.108 1573
+3DGS + Ours (L1/SSIM) 0.863 27.29 0.110 1175
 3DGS [23] (L2) 0.854 27.31 0.117 1528
@@
 In both cases, our method accelerates the convergence.
-We report results on the garden scene from the MipNeRF360 [4] dataset starting from the same initialization.
-We utilize the same loss functions in our LM optimizer as in the original 3DGS implementation, namely the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses (see Eq.
-1).
-Since LM energy terms are defined as a sum of squares, we adopt the square root formulation of these loss functions to arrive at an identical objective (see Eq.
-2).
-We compare this choice against fitting the Gaussians with only an $\mathcal{L}_{2}$ loss, that does not require taking a square root.
-Concretely, we compare the achieved quality and runtime of LM against ADAM for both the $\mathcal{L}_{2}$ loss and the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ losses.
-As can be seen in Tab.
-2, we achieve faster convergence and similar quality in both cases.
-However, the achieved quality is inferior for both LM and ADAM when only using the $\mathcal{L}_{2}$ loss.
-This highlights the importance of the $\mathcal{L}_{1}$ and $\mathcal{L}_{SSIM}$ loss functions and why we adopt them in our method as well.
-How many images per batch are necessary?
+Results on the GARDEN scene from MipNeRF360 [4].
 […]
-3DGS + Ours 5 100 0.969 33.77 0.030 242
-3DGS + Ours 5 80 0.969 33.73 0.031 233
-3DGS + Ours 5 60 0.968 33.69 0.031 223
-3DGS + Ours 5 40 0.967 33.51 0.032 212
-3DGS [20] 5 100 0.967 33.48 0.033 164
-3DGS [20] 50 100 0.968 33.57 0.032 218
-3DGS [20] 100 100 0.969 33.65 0.031 270
+100 0.969 33.77 0.030 242 32.5
+80 0.969 33.73 0.031 233 29.8
+60 0.968 33.69 0.031 223 22.6
+40 0.967 33.51 0.032 212 15.4
 Table 3: Ablation of batch-size.
-We vary the number of iterations (#I) and number of images in a batch (#B).
-Reducing #B for our method reduces runtime and consumed memory, while only slightly impacting quality.
-This demonstrates that image subsampling Sec.
-3.3.2 is compatible with LM in our task.
-Increasing #I in 3DGS [20] with ADAM is slower and achieves worse quality than Ours using LM.
-We report average results on the NeRF-Synthetic [28] dataset starting from the same initialization.
-The key hyperparameters in our model are the number of images in a batch and how many batches to choose for every LM iteration (Sec.
-3.3.2).
-This directly controls how much GPU memory our optimizer consumes and how much time we need per iteration.
-We compare different number of images in Tab.
-3 (top) on the NeRF-Synthetic [28] dataset.
-We choose this dataset because it is possible to construct the cache for all 100 training images in 33 GB of GPU memory and optimize the scene with one batch of all images.
-We compare this against our proposed subsampling scheme for different number of images in a single batch.
-Decreasing the number of images in a batch results in only slightly worse quality, but also yields faster convergence and reduces GPU memory consumption linearly down to 15GB for 40 images.
-This demonstrates that subsampling images does not negatively impact the convergence of the LM optimizer in our task.
-This motivates our usage of subsampling on the real-world datasets in our main results.
-Are we better than full-batch ADAM?
-One reason why LM requires drastically fewer iterations to converge than ADAM is the larger batch size (number of training images per iteration).
-Concretely, we require only 5 additional LM iterations after the initialization, whereas ADAM runs for another 10K iterations to converge.
-We can similarly increase the batch size in ADAM to reduce the number of iterations.
-However, as can be seen in Tab.
-3 (bottom), the achieved quality is worse for the same batch size and number of iterations.
-When running for more iterations, ADAM eventually converges to similar quality, but needs more time.
-This highlights the efficiency of our optimizer: since we solve the normal equations in Eq.
-2, one LM iteration makes a higher quality update step than ADAM which only uses the gradient direction.
-4.3 Runtime Analysis
-Figure 6: Runtime Analysis. One iteration of our LM optimizer is dominated by solving PCG and building the cache. Measured on the “garden” scene from Mip-NeRF360 [4] after densification.
-We analyze the runtime of our LM optimizer across multiple iterations in Fig.
-6.
-The runtime is dominated by solving Eq.
-3 with PCG and building the cache (Sec.
+Selecting fewer images per LM iteration reduces runtime and consumed GPU memory, while only slightly impacting quality.
+This demonstrates that image subsampling (Sec.
+3.3) is compatible with LM in our task.
+Results obtained after initialization with 3DGS [23] and with $\text{n}_{\text{b}}{=}1$.
+[…]
+3DGS [23] 10,000 1 1222 29.51
+3DGS [23] 50 75 962 29.54
+3DGS [23] 130 75 1193 29.68
++ Ours 5 75 951 29.72
+DISTWAR [12] 10,000 1 841 29.47
+DISTWAR [12] 50 75 681 29.49
+DISTWAR [12] 130 75 814 29.58
++ Ours 5 75 672 29.60
+gsplat [48] 10,000 1 919 29.52
+gsplat [48] 50 75 724 29.53
+gsplat [48] 130 75 892 29.56
++ Ours 5 75 716 29.58
+Taming-3DGS [32] 10,000 1 447 29.84
+Taming-3DGS [32] 50 75 328 29.86
+Taming-3DGS [32] 130 75 391 29.91
++ Ours 5 75 347 29.91
+Table 4: Analysis of multi-view constraints.
+We obtain higher quality update steps from our LM optimization and need fewer iterations to converge.
+Using equally many images in a batch, baselines using ADAM still require more iterations and runtime to reach similar quality.
+Results averaged on DeepBlending [19].
+Figure 6: Runtime Analysis. One iteration of our LM optimizer is dominated by solving PCG and building the cache. Measured on the GARDEN scene from Mip-NeRF360 [4] after densification.
+By replacing ADAM with our LM scheme, we accelerate the 3DGS convergence speed by 20% on average for all datasets and baselines.
+However, some drawbacks remain.
+First, our approach requires more GPU memory than baselines, due to our gradient cache (Sec.
 3.3).
-Sorting the cache, rendering all selected images, and the line search (Eq.
-4) are comparably faster.
-When running PCG, we execute the applyJ and applyJT kernels up to 8 times using the per-pixel-per-splat parallelization pattern (Algorithm 1).
-In contrast, we execute the buildCache kernel once and parallelize per-pixel, which is only marginally faster.
-This demonstrates the speed advantage obtained by using our proposed caching structure.
-We also provide a detailed profiling analysis of the kernels in the supplement.
-4.4 Limitations
-By replacing ADAM with our LM scheme, we accelerate the convergence speed by 30% on average for all datasets and baselines. However, some drawbacks remain.
-First, our approach has a higher GPU memory footprint than baselines, due to our gradient cache (Sec.
-3.3).
-Depending on the number of images and their resolution, this can make it hard to run our method on smaller GPUs.
-Following Mallick et al.
-[27], one could reduce the cache size by storing the gradients $\frac{\partial p}{\partial s}$ only for every 32nd splat along a ray and re-doing the $\alpha$-blending in these local windows.
-Second, our two-stage approach relies on ADAM to first complete the densification.
-The original 3DGS [20] densifies Gaussians up to 140 times, which is not easily transferable to the granularity of only 5 LM iterations.
-Instead, one could explore recent improvements in densification and integrate them into our method [22, 5].
+Depending on the number and resolution of images, this might require additional CPU offloading of cache parts to run our method on smaller GPUs.
+Following Mallick et al[32], one can further reduce the cache size by storing the gradients $\frac{\partial c}{\partial s}$ only for every 32nd splat along a ray and re-doing the $\alpha$-blending in these local windows.
+Second, our two-stage approach relies on ADAM for the densification.
+3DGS [23] densifies Gaussians up to 140 times, which is not easily transferable to the granularity of only 5-10 LM iterations.
+Instead, one could explore and integrate recent alternatives [25, 5, 30].
 5 Conclusion
@@
 3.2).
-We show that with our data parallization scheme we can efficiently solve the normal equations with PCG in custom CUDA kernels (Sec.
+We show that with our data parallelization scheme we can efficiently solve the normal equations with PCG in custom CUDA kernels (Sec.
 3.3).
 Employed in a two-stage approach (Sec.
-3.4), this leads to a 30% runtime acceleration compared to baselines.
-We further demonstrate that our approach is agnostic to other methods  [11, 41, 27], which further improves the optimization runtime.
-Overall, we believe that the ability of faster 3DGS reconstructions with our method will open up further research avenues and make existing 3DGS more practical across a wide range of real-world applications.
+3.4), this leads to a 20% runtime acceleration compared to baselines.
+We further demonstrate that our approach is agnostic to other methods [12, 48, 32], which further improves the optimization runtime; i.e., we can easily combine our proposed optimizer with faster 3DGS methods.
+Overall, we believe that the ability of faster 3DGS reconstructions with our method will open up further research avenues like [28] and make 3DGS more practical across a wide range of real-world applications.
 6 Acknowledgements
@@
 Appendix A More Details About CUDA Kernel Design
-We introduce the necessary CUDA kernels to calculate the PCG algorithm in Sec. 3.3.1. In this section, we provide additional implementation details.
-The necessary computations for the buildCache and applyJT steps in the PCG algorithm (see Algorithm 1) are split across three kernels.
+We introduce the necessary CUDA kernels to calculate the PCG algorithm in Sec. 3.3. In this section, we provide additional implementation details.
+A.1 Parallelization Pattern
+We implement the per-pixel-per-splat parallelization pattern in our CUDA kernels by reading subsequent entries from the gradient cache in subsequent threads.
+This makes reading cache values perfectly coalesced and therefore minimizes the overhead caused by the operation.
+The gradient cache is sorted over Gaussians, which means that subsequent entries refer to different rays (pixels) that saw this projected Gaussian (splat).
+In general, one thread handles all residuals corresponding to the respective pixel, i.e., many computations are shared across color channels and are therefore combined.
+A.2 Design Of buildCache And applyJT
+The necessary computations for the buildCache and applyJT steps in PCG (see Algorithm 1) are split across three kernels.
 This follows the original design of the 3DGS differentiable rasterizer [23].
-In both cases, we need to calculate the Jacobian-vector product of the form $\mathbf{g}=\mathbf{J}^{T}\mathbf{u}$ where $\mathbf{J}\in\mathbb{R}^{NxM}$ is the Jacobian matrix of $N$ residuals and $M$ Gaussian parameters and $\mathbf{u}\in\mathbb{R}^{N}$ is an input vector.
-The $k$-th element in the output vector is calculated as
+In both cases, we calculate the Jacobian-vector product of the form $\mathbf{g}{=}\mathbf{J}^{T}\mathbf{u}$ where $\mathbf{J}{\in}\mathbb{R}^{NxM}$ is the Jacobian matrix of $N$ residuals and $M$ Gaussian parameters and $\mathbf{u}{\in}\mathbb{R}^{N}$ is an input vector.
+The $k$-th element in the output vector is calculated as:
 $\mathbf{g}_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}\mathbf{u}_{i}=\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$ (9)
 where $\mathbf{r}_{i}$ is the $i$-th residual and $\mathbf{x}_{k}$ is the $k$-th Gaussian parameter.
+In other words, $\mathbf{g}_{k}$ is a sum over all residuals for the $k$-th Gaussian parameter.
 Following the chain-rule, it is possible to split up the gradient $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}{=}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}}$.
 We can use this to split the computation across three smaller kernels, where only the first needs to calculate the sum over all residuals: $\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\mathbf{u}_{i}$.
+This sum is the main bottleneck for the kernel implementation, since it needs to be implemented atomically (i.e., multiple threads write to the same output position).
 The other kernels then only calculate the remaining steps by parallelizing over Gaussians.
-Since the sum needs to be implemented atomically, i.e., multiple threads write to the same output position, this is the main bottleneck for the kernel implementation.
-By splitting the computation in three smaller parts, we only need to do the costly summation in the first kernel and only to intermediate attributes.
-Concretely $\mathbf{y}_{k}$ are the 2D mean, color, and opacity attributes of the $k$-th projected Gaussian.
-Since both the buildCache and applyJT kernels use the gradient cache and our proposed per-pixel-per-splat parallelization pattern, we can implement the sum by first doing a warp reduce and then only issuing one atomicAdd statement per warp.
+We slightly abuse notation and denote with $\mathbf{y}_{k}$ the 2D mean, color, and opacity attributes of the $k$-th projected Gaussian.
+That is, the first kernel sums up the partial derivatives to each of these attributes in separate vectors.
+In the following, we add the suffixes _p1, _p2, _p3 to denote the three kernels for the respective operation (where _p1 refers to the kernel that calculates the sum over residuals).
+The buildCache_p1 utilizes the original per-pixel parallelization, whereas the applyJT_p1 kernel use the gradient cache and our proposed per-pixel-per-splat parallelization pattern.
+The gradient cache is sorted over Gaussians, i.e., subsequent entries correspond to different rays of the same splat.
+This allows us to efficiently implement the sum by first performing a segmented warp reduce and then only issuing one atomicAdd statement per warp.
+A.3 Design Of applyJ And diagJTJ
 In contrast, the applyJ and diagJTJ computations cannot be split up into smaller kernels.
 Concretely, the applyJ kernel calculates $\mathbf{u}{=}\mathbf{J}\mathbf{p}$ with $\mathbf{p}{\in}\mathbb{R}^{M}$.
-The $k$-th element in the output vector is calculated as
+The $k$-th element in the output vector is calculated as:
 $\mathbf{u}_{k}=\sum_{i=0}^{M}\frac{\partial\mathbf{r}_{k}}{\partial\mathbf{x}_{i}}\mathbf{p}_{i}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{k}}{\partial\mathbf{y}_{i}}\frac{\partial\mathbf{y}_{i}}{\partial\mathbf{x}_{i}}\mathbf{p}_{i}$ (10)
-Similarly, the diagJTJ kernel calculates $\mathbf{M}=\text{diag}(\mathbf{J}^{T}\mathbf{J})\in\mathbb{R}^{M}$. The $k$-th element in the output vector is calculated as
-$\mathbf{g}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}})^{2}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}})^{2}$ (10)
+In other words, $\mathbf{u}_{k}$ is a sum over all Gaussian attributes for the $k$-th residual.
+Similarly, the diagJTJ kernel calculates $\mathbf{M}=\text{diag}(\mathbf{J}^{T}\mathbf{J})\in\mathbb{R}^{M}$.
+The $k$-th element in the output vector is calculated as:
+$\mathbf{M}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}})^{2}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{x}_{k}})^{2}$ (11)
 In both cases it is not possible to move part of the gradients outside of the sum.
-As a consequence, both the applyJ and diagJTJ are implemented as one kernel, where each thread directly calculates the complete backward path to all Gaussian attributes.
+As a consequence, both applyJ and diagJTJ are implemented as one kernel, where each thread directly calculates the final partial derivatives to all Gaussian attributes.
 This slightly increases the number of required registers and the runtime compared to the applyJT kernel (see Tab.
-4).
+5).
+The diagJTJ kernel makes use of the same segmented warp reduce as applyJT_p1 for efficiently summing up the squared partial derivatives.
+The applyJ kernel first sums up over all Gaussian attributes within each thread separately.
+Then, we only issue one atomicAdd statement for each residual per thread.
 The applyJ kernel requires the input vector $\mathbf{p}$ to be sorted per Gaussian to make reading from it coalesced.
@@
 In total, each Gaussian consists of 59 parameters: 11 for position, rotation, scaling, and opacity and 48 for all Spherical Harmonics coefficients of degree 3.
-In contrast, all other kernels require the input vector to be sorted per attribute to make reading from it coalesced.
+In contrast, all other kernels require the output vector to be sorted per attribute to make writing to it coalesced.
 That is: $\mathbf{q}{=}[x_{1}^{a},...,x_{M}^{a},...,x_{1}^{z},...,x_{M}^{z}]^{T}$.
@@
 Whenever we call the applyJ kernel, we thus first call the sortX kernel that restructures $\mathbf{q}$ to the layout of $\mathbf{p}$.
-Appendix B Derivation of Weights for Subsampling
-We sample batches of fewer images to decrease the size of the gradient cache (see Sec.
-3.3.2).
-To combine the update vectors from multiple batches, we calculate the weighted mean, as detailed in Eq.
-7.
-This weighted mean approximates the true solution without any image subsampling.
-That is, to obtain the update vector, we have to solve the normal equations, as detailed in Eq.
-3.
-When subsampling images, we split the number of total residuals $M$ into smaller chunks.
-Let’s consider the case of two chunks, labeled as $1$, $2$.
-The normal equations (without subsampling) can then be written as:
+A.4 Precomputation Of Residual-To-Pixel Weights
+We adopt the square root formulation of the residuals in our energy formulation (see Eq.
+3).
+We efficiently precompute the contribution of the square root to the partial derivatives of Eq.
+9, Eq.
+10, and Eq.
+11.
+In the following, we divide the partial derivatives from the $i$-th residual to the $k$-th Gaussian attribute into two stages:
+$\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}=\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ (12)
+where $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$ goes from the residual to the rendered pixel color and $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ from the pixel color to the Gaussian attribute.
+Since we adopt the L1 and SSIM loss terms and take their square root, the terms $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$ need to be calculated accordingly.
+In contrast, when using the L2 loss they take a trivial form of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}{=}1$.
+In the following, we show that we can simplify the calculation of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}$ in the kernels by precomputing $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$.
+The $k$-th element of $\mathbf{g}$ is calculated as:
+$\mathbf{g}_{k}=(\mathbf{J}^{T}\mathbf{J}\mathbf{p})_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}\sum_{j=0}^{M}\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}$ (13)
+By substituting Eq. 12 into Eq. 13, we factor out $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}}$:
+$\mathbf{g}_{k}=\sum_{i=0}^{N}(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}$ (14)
+The terms $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ are identical for a residual of the same pixel and color channel that corresponds to either the L1 or SSIM loss terms, respectively.
+To avoid computing $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ twice (and therefore doubling the grid size of all kernels), we instead sum up the contribution of both loss terms:
+$(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}=(\frac{\partial\mathbf{r}_{i}^{1}}{\partial\mathbf{c}_{i}})^{2}+(\frac{\partial\mathbf{r}_{i}^{2}}{\partial\mathbf{c}_{i}})^{2}$ (15)
+where $\mathbf{r}_{i}^{1}$ corresponds to the $i$-th L1 residual and $\mathbf{r}_{i}^{2}$ to the $i$-th SSIM residual.
+Additionally, we implement the multiplication with $(\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{c}_{i}})^{2}$ in Eq.
+14 as elementwise vector product (denoted by $\odot$) of $\mathbf{\hat{u}}{=}[\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{0}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}...\sum_{j=0}^{M}\frac{\partial\mathbf{c}_{N}}{\partial\mathbf{x}_{j}}\mathbf{p}_{j}]$ and $\mathbf{\nabla r}{=}[(\frac{\partial\mathbf{r}_{0}}{\partial\mathbf{c}_{0}})^{2}...(\frac{\partial\mathbf{r}_{N}}{\partial\mathbf{c}_{N}})^{2}]$:
+$\mathbf{g}_{k}=\sum_{i=0}^{N}\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}(\mathbf{\hat{u}\odot\nabla r})_{i}$ (16)
+This avoids additional uncoalesced global memory reads to $\mathbf{\nabla r}$ in the CUDA kernels.
+Instead, we calculate $\mathbf{\hat{u}\odot\nabla r}$ in a separate operation after the applyJ kernel and before applyJT.
+This also simplifies the kernels, since they now only need to compute $\frac{\partial\mathbf{c}_{i}}{\partial\mathbf{x}_{k}}$ instead of $\frac{\partial\mathbf{r}_{i}}{\partial\mathbf{x}_{k}}$.
+Therefore, the only runtime overhead of using the L1 and SSIM residual terms over the L2 residuals is the computation of $\mathbf{\nabla r}$.
+However, this can be efficiently computed using backpropagation (autograd) and is therefore not a bottleneck.
+Appendix B Derivation Of Image Subsampling Weights
+We subsample batches of images to decrease the size of the gradient cache (see Sec.
+3.3).
+To combine the update vectors from multiple batches, we calculate their weighted mean, as detailed in Eq.
+8.
+This weighted mean approximates the “true” solution to the normal equations (Eq.
+4), that does not rely on any image subsampling (and instead uses all available training images).
+When subsampling images, we split the number of total residuals into smaller chunks.
+In the following, we consider the case of two chunks (labeled as 1 and 2), but the same applies to any number of chunks.
+We re-write the normal equations (without subsampling) using the chunk notation as:
 $\begin{bmatrix}\mathbf{J}^{T}_{1}&\mathbf{J}^{T}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{J}_{1}\\ \mathbf{J}_{2}\end{bmatrix}\Delta=\begin{bmatrix}\mathbf{J}^{T}_{1}&\mathbf{J}^{T}_{2}\end{bmatrix}\begin{bmatrix}\mathbf{F}_{1}(\mathbf{x})\\ \mathbf{F}_{2}(\mathbf{x})\end{bmatrix}$ (17)
-where we drop the additional LM regularization term for clarity and divide the Jacobian and residual vector into separate matrices according to the chunks.
+where we drop the additional LM regularization term for clarity and divide the Jacobian and residual vector into separate matrices/vectors according to the chunks.
 The solution to the normal equations is obtained by:
 $\Delta=(\mathbf{J}^{T}_{1}\mathbf{J}_{1}+\mathbf{J}^{T}_{2}\mathbf{J}_{2})^{-1}(\mathbf{J}^{T}_{1}\mathbf{F}_{1}(\mathbf{x})+\mathbf{J}^{T}_{2}\mathbf{F}_{2}(\mathbf{x}))$ (18)
-In contrast, when we subsample images, we obtain two separate solutions as:
+In contrast, when we subsample images, we solve the normal equations separately and obtain two solutions:
 $\Delta_{1}=(\mathbf{J}^{T}_{1}\mathbf{J}_{1})^{-1}\mathbf{J}^{T}_{1}\mathbf{F}_{1}(\mathbf{x})$ (19)
@@
 This directly leads to the weighted mean that we employ in Eq.
-7.
-Appendix C Detailed Analysis of Runtime
-Kernel Duration (ms) $\downarrow$ Compute Throughput (%$\uparrow$ Memory Throughput (%$\uparrow$ Register Count $\downarrow$
+8.
+Appendix C Detailed Runtime Analysis
+[…]
 buildCache_p1 31.32 78.56 78.56 64
@@
 We provide results measured on a RTX3090 GPU for building/resorting the gradient cache and running one PCG iteration on the MipNerf360 [4] “garden” scene with a batch size of one image.
-We provide additional analysis of the CUDA kernels by running the Nsight Compute [^1] https://developer.nvidia.com/nsight-compute profiler.
+We provide additional analysis of the CUDA kernels by running the NVIDIA Nsight Compute profiler.
 We provide results in Tab.
@@
 We add the suffixes _p1, _p2, _p3 to signal the three kernels that we use to implement the respective operation (see Appendix A).
+Comparing the runtime of the buildCache and applyJT kernels reveals the advantage of our proposed per-pixel-per-splat parallelization pattern.
+Both compute the identical Jacobian-vector product, but the buildCache kernel relies on the per-pixel parallelization pattern of the original 3DGS rasterizer [23].
+However, we compute the result 4.84.8x faster using the gradient cache in the applyJT kernel.
+We also note that the compute and memory throughput as well as the register count of both kernels are roughly similar.
+This signals that our kernel implementation is equally efficient, i.e., there are no inherent drawbacks using our proposed GPU parallelization scheme.
 Appendix D Results Per Scene
@@
 + Ours truck 0.876 25.36 0.151 747
-Table 5: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
+Table 6: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the 3DGS [23] baseline.
 Method Scene MipNeRF-360 [4]
@@
 + Ours truck 0.877 25.28 0.150 635
-Table 6: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
+Table 7: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the DISTWAR [12] baseline.
 Method Scene MipNeRF-360 [4]
@@
 + Ours truck 0.877 25.28 0.150 447
-Table 7: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
+Table 8: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics against the gsplat [48] baseline.
 Method Scene MipNeRF-360 [4]
 […]
-Taming-3DGS [27] treehill 0.630 22.61 0.340 638
-+ Ours treehill 0.635 22.53 0.332 510
-Taming-3DGS [27] counter 0.904 28.96 0.205 583
-+ Ours counter 0.904 28.91 0.205 553
-Taming-3DGS [27] stump 0.764 26.51 0.225 728
-+ Ours stump 0.773 26.70 0.217 579
-Taming-3DGS [27] bonsai 0.938 32.16 0.208 508
-+ Ours bonsai 0.938 31.92 0.208 468
-Taming-3DGS [27] bicycle 0.756 25.18 0.224 882
-+ Ours bicycle 0.764 25.26 0.218 702
-Taming-3DGS [27] kitchen 0.924 31.13 0.129 742
-+ Ours kitchen 0.924 31.14 0.128 696
-Taming-3DGS [27] flowers 0.594 21.42 0.351 620
-+ Ours flowers 0.598 21.48 0.348 507
-Taming-3DGS [27] room 0.916 31.64 0.224 561
-+ Ours room 0.916 31.39 0.224 529
-Taming-3DGS [27] garden 0.861 27.36 0.113 895
-+ Ours garden 0.861 27.32 0.112 748
+Taming [32] treehill 0.625 23.00 0.385 479 0.78
++ Ours treehill 0.623 23.03 0.332 381 0.78
+Taming [32] counter 0.896 28.59 0.223 646 0.31
++ Ours counter 0.894 28.51 0.205 537 0.31
+Taming [32] stump 0.734 25.96 0.292 405 0.48
++ Ours stump 0.733 25.98 0.217 315 0.48
+Taming [32] bonsai 0.934 31.73 0.221 634 0.41
++ Ours bonsai 0.932 31.64 0.208 504 0.41
+Taming [32] bicycle 0.716 24.78 0.295 485 0.81
++ Ours bicycle 0.709 24.75 0.218 376 0.81
+Taming [32] kitchen 0.918 30.85 0.141 722 0.48
++ Ours kitchen 0.918 30.84 0.128 597 0.48
+Taming [32] flowers 0.554 21.00 0.407 465 0.57
++ Ours flowers 0.549 20.99 0.348 365 0.57
+Taming [32] room 0.906 31.12 0.251 621 0.22
++ Ours room 0.906 31.21 0.224 521 0.22
+Taming [32] garden 0.852 27.24 0.128 638 1.90
++ Ours garden 0.852 27.25 0.112 483 1.90
 Method Scene Deep Blending [19]
 […]
-Taming-3DGS [27] playroom 0.903 29.99 0.245 502
-+ Ours playroom 0.904 30.23 0.247 453
-Taming-3DGS [27] drjohnson 0.902 29.43 0.240 678
-+ Ours drjohnson 0.901 29.21 0.249 627
+Taming [32] playroom 0.900 30.29 0.278 419 0.18
++ Ours playroom 0.902 30.41 0.280 324 0.18
+Taming [32] drjohnson 0.899 29.38 0.269 475 0.40
++ Ours drjohnson 0.899 29.40 0.271 370 0.40
 Method Scene Tanks & Temples [27]
 […]
-Taming-3DGS [27] train 0.815 22.18 0.205 411
-+ Ours train 0.809 22.06 0.217 349
-Taming-3DGS [27] truck 0.879 25.40 0.146 514
-+ Ours truck 0.879 25.36 0.150 436
-Table 8: Quantitative comparison of our method and baselines. We show the per-scene breakdown of all metrics.
+Taming [32] train 0.815 22.18 0.205 411 0.36
++ Ours train 0.802 22.28 0.241 328 0.36
+Taming [32] truck 0.879 25.40 0.146 514 0.27
++ Ours truck 0.862 25.16 0.178 292 0.27
+Table 9: Quantitative comparison of our method and baselines.
+We show the per-scene breakdown of all metrics against the Taming-3DGS [32] baseline.
+Additionally, we include the number of Gaussians in millions (#G (M)) that we obtained using the default hyperparameters for “budgeting”.
 We provide a per-scene breakdown of our main quantitative results against all baselines on all datasets.
 The comparisons against 3DGS [23] are in Tab.
-5.
+6.
 The comparisons against DISTWAR [12] are in Tab.
-6.
+7.
 The comparisons against gsplat [48] are in Tab.
-7.
+8.
 The comparisons against Taming-3DGS [32] are in Tab.
-8.
+9.
+Our method shows consistent acceleration of the optimization runtime on all scenes, while achieving the same quality.